### v1

# SynthCoder: A Synthetical Strategy to Tune LLMs for Code Completion

Abstract
Code completion is a prominent application of Large Language Models (LLMs) in software engineering. Due to the near real-time response requirements of this task, base models with small to medium-sized parameters are typically employed, supplemented by various optimization and post-training techniques. However, these optimization methods often have trade-offs, leading to a seesaw effect where performance improvements on certain datasets or metrics are accompanied by degradations on others – sometimes even falling below the baseline model’s performance. This paper proposes SynthCoder, a model that integrates leading industry practices to achieve state-of-the-art performance on the Fill-in-the-Middle (FIM) code completion task. In specific, we first construct a diverse dataset by combining Abstract Syntax Tree (AST) node extraction with heuristics that simulate developer behavior. Then we enrich our training corpus with cross-file contextual information using the BM25 algorithm and call graphs, enhancing the model’s ability to perform code completion in both file-level and repository-level scenarios. As the last step, we employ a two-stage training process using the Seed-Coder-8B-Base as the base model. First, we fine-tune the model using Curriculum Learning technology. Following this, we perform alignment using Direct Preference Optimization (DPO) with preference pairs generated through Rejection Sampling. Experimental results demonstrate that our final model excels on mainstream repository-level code completion benchmarks, including aiXcoder, ExecRepoBench, CrossCodeEval, and CoLT. Furthermore, our carefully curated training set effectively mitigates the model’s tendency to just repeat existing code, a common issue existing in various code completion models.
Index Terms:
LLM, code completion, fill in the middle
I Introduction
Code completion aims to accelerate software development by predicting a developer’s subsequent coding intentions and providing syntax hints and code snippets. Early code completion approaches were primarily based on predefined rules and traditional machine learning methods. They could index the context and offer suggestions as developers typed keywords and identifiers, significantly alleviating the burden of memorizing precise syntax, function names, or Application Programming Interface (API) details [1, 2]. As software systems have grown in complexity, so too have developer expectations for code completion. Traditional methods, however, have struggled to meet these more demanding requirements. In recent years, Large Language Models (LLMs) have brought transformative changes to the field with their powerful capabilities in text understanding, reasoning, and generation [3]. A multitude of products integrating LLMs with code completion have emerged in the market. Examples include the GitHub Copilot  [^1] https://github.com/features/copilot plugin, which is compatible with various IDEs, and the Cursor  [^2] https://www.cursor.com IDE, which deeply integrates LLMs into various features of the development tool. These products are often built upon proprietary, cloud-based commercial models from providers like OpenAI and Anthropic, and have achieved remarkable success in both performance and reputation. One study demonstrated that using GitHub Copilot can increase programming speed by 55.8% compared to a control group [4]. Furthermore, a survey revealed that developers experience reduced cognitive load and higher job satisfaction after adopting AI-powered coding tools  [^3] https://github.blog/news-insights/research/research-quantifying-github-copilots-impact-on-developer-productivity-and-happiness. Concurrently, the industry are also developing open-source alternatives, such as Tabby  [^4] https://github.com/TabbyML/tabby.
Unlike conversational models, the input for code completion usually consists of the code surrounding the target location and other relevant context. The desired output is strictly confined to the code for that location, which requires the model to understand the input and precisely constrain its output. The Fill-in-the-Middle (FIM) paradigm, proposed by OpenAI [5], established a specialized input-output format for code completion that has since been widely adopted. In the domain of code completion models, open-source weights from models like CodeLlama [6] and fully open-sourced training data from models like StarCoder [7] have provided valuable references for both academia and industry. Typically, the Base version of a model series is designed for FIM code completion tasks, whereas the Instruct version is more adept at conversational tasks [8, 9]. However, directly applying these models to code completion often yields unsatisfactory performance across different scenarios. Fortunately, base models are able to serve as an excellent starting point for post-training. We can construct training dataset that better reflects real-world scenarios to further train these models and enhance their FIM capabilities. A noteworthy point is that since in-line code completion is usually triggered following the developer’s cursor, response speed is a critical factor  [^5] https://sourcegraph.com/blog/the-lifecycle-of-a-code-ai-completion. Consequently, many FIM code completion models have fewer than 10B active parameters [10, 11, 9], or include versions within this range [6, 8, 7]. Therefore, the scope of this paper is thus also limited to models with fewer than 10B active parameters.
The main contributions of this paper are as follows:
  1. 1.
    We propose a comprehensive methodology for code completion models that organically integrates and improves upon the data construction techniques, and training methods from several existing works [12, 13, 10], resulting in a new state-of-the-art (SOTA) model for FIM code completion.
  2. 2.
    We empirically evaluate several mainstream code completion models and discuss the gap between model capabilities and their successful application in completion tools, specifically addressing the issue of models erroneously repeating one or more lines of existing context.
The structure of this paper is as follows. Section III introduces the background and motivation of our research, and Section II reviews related work on applying LLMs to domains such as code completion. Section III provides a detailed account of our entire process, from data acquisition (Section III-A) and data processing (Section III-B) to training details (Section III-C). To validate the effectiveness of our method, in Section IV, we conduct a comprehensive evaluation of all baseline models and SynthCoder on file-level and repository-level code completion tasks, followed by an in-depth analysis of the results. Based on these experimental findings, Section V discusses the gap between code completion performance on test sets versus in practical applications. Section VI addresses the limitations of the current work. Finally, Section VII concludes the paper and suggests directions for future work.
II Related Work
This section describes relevant work from the perspectives of code completion and adopting LLMs for code completion tasks.
II-A Fill in the Middle Code Completion
Large Language Models (LLMs) for code have demonstrated significant potential in code completion tasks. Early works such as StarCoder and CodeLlama laid a foundational basis for this area [7, 6]. However, constructing high-quality training data remains a challenge, particularly in ensuring that the completion target is meaningful in relation to its surrounding context (the prefix and suffix) and covers complex, real-world scenarios. Research in this domain has focused on structuring completion data to enhance model performance in authentic development contexts.
Similarly leveraging code’s structural information, the ExecRepoBench team introduced a multi-level code completion method that also relies on ASTs to build training data. Their approach emphasizes enhancing the model’s capabilities in practical coding scenarios that involve complex, multi-file interdependencies [12]. This reflects a trend in code completion research, which is progressively moving from the single-file level to more complex repository-level scenarios that include cross-file dependencies.
II-B Repository Level Code Completion
Understanding the complex interdependencies among multiple files is crucial for handling modern software projects, a task known as Repository-Level Code Completion. The Qwen2.5-Coder-Instruct-C model utilizes structural information, such as Abstract Syntax Trees (ASTs), to enhance its functionality in practical coding scenarios involving complex, multi-file dependencies, and it introduced the ExecRepoEval benchmark [12]. Similarly, CrossCodeEval employs static analysis to ensure that completion targets require cross-file information (e.g., resolving undefined API calls), thereby accurately measuring a model’s capabilities in realistic multi-file collaboration scenarios [14].
In addition to enhancing the intrinsic capabilities of models, Retrieval-Augmented Generation (RAG) has demonstrated significant potential for improving repository-level code completion by enriching the context provided to the model. The RepoCoder framework proposed a two-stage interactive retrieval process: the first stage locates relevant files, and the second extracts fine-grained code snippets, effectively addressing the information fragmentation problem of single-pass retrieval [15]. Taking this further, RepoFormer innovatively introduced a self-triggered retrieval mechanism, training the model to generate special tokens (e.g., <cc>) to autonomously decide when cross-file retrieval is necessary [16]. Structured retrieval strategies have also seen breakthroughs: GraphCoder constructs a Code Context Graph (CCG) that integrates relationships like control flow and data dependencies, optimizing retrieval results using graph edit distance [17]. Similarly, RepoHyper utilizes an extended Repository-level Semantic Graph (RSG) and a link prediction-based retrieval strategy to significantly enhance cross-file code completion capabilities [18]. In terms of architectural innovation, the Code Graph Model (CGM) represents a repository as a heterogeneous graph (containing seven types of nodes and five types of edges) and integrates this structural information into the LLM’s attention layers via a graph-aware attention mask [19].
II-C Post-training LLMs for Coding
Code generation is a significant application of LLMs. The core technology is founded on the Transformer architecture, which replaced Recurrent Neural Networks (RNNs) [20] and Convolutional Neural Networks (CNNs) [21]. By leveraging a self-attention mechanism, Transformers effectively model long-range dependencies and exhibit excellent scaling stability, laying the groundwork for learning from code corpora [22].
The decoder-only architecture has become the mainstream choice for code LLMs due to its advantages in text generation, including code. GPT-3 (175B) [23] demonstrated general-purpose capabilities (including code generation) without fine-tuning, and its emergent abilities [24] greatly spurred exploration in the field. Subsequently, OpenAI trained the CodeX model on a vast amount of open-source code (e.g., based on GPT-3.5), revealing for the first time the broad prospects of LLMs in software engineering.
The open-source community has since produced powerful base models, such as Llama/Llama2 [25, 26], the Qwen series, and the DeepSeek series, as well as code-specific models like DeepSeek-Coder [27], Qwen2.5-Coder [8], and Seed-Coder [9]. The base model used for fine-tuning in this paper is from the Seed-Coder series. These models have achieved outstanding results on code benchmarks, with some metrics even surpassing those of larger, closed-source models.
In terms of expanding model capabilities, the o1 model launched by OpenAI in 2024 opened a new dimension: Inference Time Scaling [28]. These models are also known as Reasoning LLMs. Shortly after, the DeepSeek-R1 [29] model was open-sourced, presenting a method for training reasoning models using reinforcement learning that achieved complex task-solving abilities comparable to o1. Although reasoning models are constrained in low-latency tasks like code completion, they are becoming indispensable for more systematic software engineering tasks such as code repair, repository-level code understanding, and complex code generation.
Another critical path to enhancing code LLM performance is the construction of high-quality, domain-specific training data. This is particularly challenging for code completion (the FIM task, as described in Section II-A) and code instruction generation. For the latter, the primary challenge is to increase data diversity and difficulty. Early work like Code Alpaca generated instruction-output pairs from seed tasks using ChatGPT [30], but its data diversity was limited. To address this, subsequent research proposed innovative methods: Evol-Instruct [31] evolutionarily rewrites simple instructions, while OSS-Instruct [32] innovatively uses open-source code as seeds to prompt an LLM to generate relevant problems and solutions.
In summary, current research on code Large Language Models focuses on three key directions: (1) Constructing high-quality Fill in the Middle (FIM) training data by leveraging structural information like Abstract Syntax Trees (ASTs) to enhance the model’s code completion capabilities in both single-file and complex cross-file dependency scenarios [11, 12]; (2) Advancing the code completion task from the single-file level to the Repository Level, which requires an understanding of the overall project structure and cross-file dependencies. Relevant research encompasses innovative paradigms such as retrieval augmentation [15, 16], graph structure modeling [17, 15], and the integration of static analysis [33, 19]; (3) Continuously improving model performance on complex tasks like code generation, understanding, and repair by combining advanced post-training techniques – especially diverse instruction data synthesis methods [31, 32] – with innovative model architectures and optimization strategies [28]. Collectively, these research streams lay the foundation for building more powerful and practical intelligent code assistants. Unfortunately, due to inherent constraints in their internal mechanisms, these different approaches cannot be simply combined to boost the performance of code completion models [34].
III Methodology
The details in training SynthCoder will be elaborated in this section.
III-A Data Retrieval
To build a high-quality and multi-dimensional training corpus for code completion, we constructed data for different application scenarios on two levels. First, we utilized The-Stack-V2 [35] open-source repository dataset to create Fill-in-the-Middle (FIM) code completion samples that include both intra-file and cross-file contexts by masking abstract syntax tree nodes. Second, based on the pre-training corpus from the OpenCoder project, we generated samples for completing function bodies based on function signatures and docstrings. Finally, to improve training efficiency and model performance, we applied a unified quality filtering process to all data, which combined a heuristic rule pipeline with sample selection based on model perplexity.
III-A1 Files from Open-Source Repositories
To construct a code completion dataset with both intra-file and cross-file contexts, with considerations of data quality and the costs of data acquisition, we extracted files in Java, Python, C++, Go, JavaScript, and TypeScript from repositories with 256 or more stars in the the-stack-v2-train-smol-ids  [^6] https://huggingface.co/datasets/bigcode/the-stack-v2-train-smol-ids, a subset of The-Stack-V2, which served as the basis for our AST node masking and cross-file context construction.
III-A2 Snippets of Standalone Functions
For the scenario of completing function bodies based on function signatures and docstrings, we used the pre-training annealing data, opc-annealing-corpus  [^7] https://huggingface.co/datasets/OpenCoder-LLM/opc-annealing-corpus/, from the OpenCoder project to construct function completion samples. This dataset contains over ten million lines of code snippets synthesized by GPT-4. Specifically, we extracted function implementations from complete code files to serve as the code snippets to be completed, while the function signature, docstring, and other external code related to the target function collectively formed the context. For Java, Python, C++, Go, JavaScript, and TypeScript, we used data from the synthetic_code_snippet subset. As the Go language was not included in this subset, we used the synthetic_qa subset for extraction instead. Since functions in this dataset may lack docstrings, we used the DeepSeek-V3-0324  [^8] https://huggingface.co/deepseek-ai/DeepSeek-V3-0324 model to synthesize corresponding docstrings for the extracted functions.
III-A3 Filtering Based on Data Quality
To improve training efficiency and enhance the final model’s performance, all data from our sources underwent a quality filtering process.
OpenCoder Pipeline
OpenCoder Data Filtering Pipeline is a heuristic filtering framework for large-scale code pre-training corpora, which considers the unique characteristics of different programming languages and includes over 100 filtering rules. The guiding principle of this pipeline is to remove harmful data while ensuring that the overall distribution of the dataset is not significantly affected [13]. In this paper, we adopt this pipeline  [^9] https://github.com/OpenCoder-llm/opc_data_filtering to filter the raw code data based on heuristic rules, thereby obtaining high-quality yet diverse code, which in turn improves the overall quality of the subsequently synthesized code completion data.
Perplexity Filtering
Perplexity (PPL) measures the certainty of a model’s predictions for a given sample. A model that outputs a low perplexity for a sample is better at predicting it [36]. For a sequence $x=(x_{1},x_{2},...,x_{N})$ of length $N$, where the model’s predicted conditional probability is $P(x_{t}|x_{<t})$, the perplexity $\text{PPL}(x)$ is as follows:
$\text{PPL}(x)=\exp(-\frac{1}{N}\sum_{t=1}^{N}\log P(x_{t}|x_{<t}))$
In this work, we compute the perplexity of the code completion and code infilling data on the LLM to be trained. The distribution of the perplexity values generally approximates a log-normal distribution. For the post-training process, data with very low perplexity represents knowledge the LLM has already mastered and can output correctly without further training. Conversely, data with very high perplexity deviates significantly from the distribution of the LLM’s base data, and this subset tends to have a higher proportion of erroneous data. We remove a certain percentage of the data with the lowest and highest perplexity to improve the overall quality of the training set [37, 38]. For the infilling data, we ultimately remove data outside of two standard deviations (2 $\sigma$) of the fitted log-normal distribution. For the completion data, we discard the lowest and highest 5% of the data.
III-B Data Construction
The core of our data construction lies in deeply simulating the real-world coding habits and contextual needs of developers. Our method encompasses three key aspects: (1) For constructing the completion target, we combine fine-grained node extraction based on Abstract Syntax Trees (ASTs) with heuristic strategies that mimic user behavior. (2) For building the completion context, we introduce critical cross-file information using code similarity (BM25) and code dependency graphs (Call Graphs). (3) For optimizing data quality, we employ rejection sampling and a multi-stage filtering process to generate high-quality preference pairs for alignment training, aiming to enhance the model’s FIM capabilities
and suppress context repetition.
Through this multi-dimensional construction strategy, we aim to generate high-quality training data that more closely resembles real-world application scenarios.
III-B1 Inner File Context
Today, AI-driven coding tools like GitHub Copilot  [^10] https://github.com/features/copilot and Cursor  [^11] https://www.cursor.com have been widely adopted by the developer community. Because developers have certain inherent habits in their thinking and coding processes – habits which are in turn influenced and reshaped by these AI tools – the locations where they expect to trigger code completion tend to follow specific patterns. Based on this core idea, our work strives to construct code completion data that aligns as closely as possible with developer habits. We have designed multiple strategies for this purpose, which can be broadly categorized into two main types: a strategy based on Abstract Syntax Tree (AST) nodes (60% of the data) and a heuristic strategy based on user habits (40%).
Strategies Based on AST Nodes
Developers tend to think and code in terms of syntactic nodes that carry specific meaning, rather than handling characters or lines of code in isolation. This observation stems from the top-down, divide-and-conquer strategy prevalent in software development, where developers decompose functionality into multiple, relatively independent code snippets. This approach significantly reduces the cognitive load of developing complex software. After breaking down complex features, developers can concentrate on implementing smaller, logically simpler code segments. This method of progressively decomposing a complex system naturally aligns with how parsing tools break code down into a tree-like structure.
In this paper, we use the Tree-sitter  [^12] https://tree-sitter.github.io/tree-sitter parsing tool to parse code files into abstract syntax trees. Based on a variety of node extraction strategies, we select target nodes as the code snippets to be completed and extract their corresponding prefix and suffix contexts. Compared to previous related work [39, 11, 12], we categorize syntactically or functionally similar nodes together, thereby creating 13 more fine-grained data synthesis strategies. These are based on: classes and interfaces; methods; formal and actual parameters; imported modules; expressions; assignment operations and variable declarations; comments; conditional statements; loop statements; return statements; other common statements; annotations or decorators; and core concurrent programming statements (only in Go).
Strategies Based on User Behavior Patterns
By summarizing developer habits when using code completion tools, this paper introduces additional construction strategies tailored to these behaviors. Specifically, this includes the following three types of data.
  • Code Completion Following a Comment: When a model cannot directly understand a developer’s intent, the developer may manually write a comment describing the functionality of the code block to be written, then trigger code completion on the next line. To simulate this, we use Tree-sitter to detect full-line comment nodes in code files and treat the first code block immediately following the comment as the completion target.
  • Code Completion Within Parentheses: Paired parentheses often enclose relatively independent code snippets. In code editors, typing an opening parenthesis automatically generates the closing one and places the cursor in between. We sample data to replicate this common scenario.
  • Completion of Incomplete Single Lines: When a developer pauses typing and there is no code to the right of the cursor, code completion is often triggered to finish the current line. We designed two strategies for this scenario.
    • Completion based on a random position: we randomly select a point that is not at the beginning or end of a line and designate the rest of the line as the completion target.
    • Completion based on special symbols: we compile a list of specific keywords and special characters for each language, search for these tokens during construction, and use the remainder of the line as the target.
To enhance the LLM’s generalization capabilities, we also randomly select single lines and consecutive multi-line blocks of code as completion targets, which accounts for approximately 10% of the total data volume.
III-B2 Cross File Context
Effective repository-level code completion requires Large Language Models (LLMs) to leverage cross-file context, moving beyond the limitations of single-file analysis to better reflect real-world development environments. To this end, we introduce a training data synthesis strategy that incorporates in-project context along two dimensions: code similarity and code dependency. Specifically, we employ the BM25 algorithm to retrieve textually similar code and leverage code call graphs to identify functional dependencies. By enriching the training data with this relevant cross-file information, we enhance the model’s ability to perform more accurate and context-aware code completion.
Retrieving Similar Code Snippets
We use BM25 algorithm [40] to retrieve code snippets similar to the target completion, providing the LLM with inspiration for analogical learning in terms of implementation logic and local code structure. Before computing text similarity, we first segment the code within the project into chunks based on blank lines, ensuring that each snippet is fewer than 20 lines long while remaining relatively self-contained [41]. The query consists of the code to be completed along with some of its adjacent context. We then compute the BM25 similarity score between this query and all code chunks from other files within the project, selecting the highest-scoring snippets (sum up within a maximum total length after tokenization) to serve as cross-file context.
Dependency Related Code Snippets
Code import statements directly indicate which external APIs and modules can be called within a file. Retrieving and providing such code information from the project allows the LLM to better decide when to use imported modules and how to use them more correctly.
This paper uses the open-source Repo-specific Semantic Graph (RSSG) call graph construction module from the RepoFuse project [42] to directly build call relationships between different files from the project structures of open-source repositories collected from The-Stack-V2 dataset. This process allows for the acquisition of each file’s imported dependencies within the project. Since the total token length of multiple complete files is likely to exceed the context length limit of the training data, this paper uses Tree-sitter to create extraction scripts tailored to the specifics of each programming language. These scripts extract class definitions, method signatures, functional description comments, and other key information from the called code files, thereby significantly reducing the context length while retaining essential information.
III-B3 Rejection Sampling
To further optimize the model’s FIM capabilities, this paper introduces paired preference alignment training. The core of this training is to construct a high-quality preference dataset, for which a negative sample generation and filtering pipeline was designed. The main steps are as follows:
Rejection Sampling Strategies
Based on a cross-file code dataset constructed using the BM25 algorithm and call relationship graphs, we use the target model intended for alignment on each sample. With the sampling temperature set to 1.0, we generate 10 candidate negative samples via rejection sampling [10].
Multi-stage Filtering Strategies
To ensure the quality and diversity of the negative samples, we apply a series of filters to the candidates. First, validity filtering is performed, which includes deduplication, removal of null values, and elimination of samples that are identical to or contain the ground truth. Subsequently, similarity filtering based on BLEU scores is used to remove negative samples that are too similar to the ground truth.
Contextual Repetition Suppression
After the screening process, we retain up to 3 negative samples for each original data point, which constitute the paired dataset for preference alignment training. At the same time, to suppress the model’s tendency to simply repeat the context, we randomly sampled 10% of the data from rejection sampling where the expected completion does not start with the first line of the suffix. For this subset, we constructed suffix-repetition suppression pairs where the positive example is the ground-truth completion and the negative example is the first line of the suffix code. This data was then added to the synthesized dataset. Similarly, we also added 1% of prefix-repetition suppression data.
III-C Model Post-training
Using the data synthesized with the methods described above, we conducted two stages of post-training on the Seed-Coder-8B-Base model [9]: fine-tuning and alignment. Seed-Coder-8B-Base is an 8B-parameter open-source large language model for code developed by the ByteDance Seed team. It serves as the base model for the Seed-Coder series and is designed to provide a powerful starting point for various downstream tasks, including code generation and reasoning. Thus, we use Seed-Coder-8B-Base as our based model for code completetion task.
III-C1 Finetune Stage
In the fine-tuning stage, we adopted the curriculum learning method from [43].
Curriculum Learning
The data preprocessing for curriculum learning fine-tuning is primarily divided into three steps:
  • Step 1: Code Parsing.
    After initial filtering, we use Tree-sitter to parse code files and generate Abstract Syntax Trees (ASTs).
  • Step 2: Complexity Sorting.
    We measure complexity by the number of identifier nodes in the AST and sort them in descending order. Symbols are extracted by recursively traversing the AST.
  • Step 3: Sample Selection.
    We select the top $k\%$ (k=30) most complex fragments to serve as FIM and completion training samples, respectively, forming the final training dataset.
Training Parameters and Cost
In the curriculum learning fine-tuning stage, we trained the model for one epoch using the AdamW optimizer, with a learning rate set to 3e-6 combined with a 50-step warmup. The training data selected was the top 30% (k=0.3) of the original dataset sorted by complexity, and the input sequences were uniformly truncated to within 8000 tokens. The entire process occupied 4 H800 80G PCIE GPUs, with a global batch size of 64 (micro-batch size of 1). Strategies such as DeepSpeed ZeRO-1, Flash Attention, and gradient checkpointing activation re-computation were used to improve training efficiency.
III-C2 Alignment Stage
To make the code generation model more accurately align with the intentions and preferences of human developers, this paper uses the Direct Preference Optimization (DPO) algorithm [44] to align the model during the preference alignment stage.
DPO Algorithm
Compared to the traditional RLHF framework, which relies on an explicit reward model and reinforcement learning, DPO provides a more stable and efficient end-to-end optimization path. The core idea of DPO is to directly transform the optimization objective on a preference dataset into a maximum likelihood estimation problem for the language model. Given a preference dataset $\mathcal{D}=\{(x^{i},y_{w}^{i},y_{l}^{i})\}$ consisting of a prompt $x$, a winning code completion $y_{w}$, and a losing code completion $y_{l}$, the goal of DPO is to fine-tune the policy model $\pi_{\theta}$ to better rank human preferences. The DPO algorithm optimizes the model by minimizing the following loss function:
$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\Bigg{[}$
$\quad\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\Bigg{]}$
Training Parameters and Cost
In the preference alignment stage using the DPO algorithm, the DPO parameter $\beta$ was set to 0.9, with an equivalent batch size of 64. The model was trained for one epoch with a maximum learning rate of 5e-5, a linear learning rate warmup of 48 steps, and a cosine learning rate decay scheduler. The maximum length for training data was 8192. For LoRA, the rank and alpha were set to 32 and 64, respectively, and the LoRA training modules included various modules within Self-Attention and the feed-forward networks. The training was conducted on 4 NVIDIA H800 80G PCIE GPUs using DeepSpeed ZeRO-2 and a CPU offloading strategy. The changes in Loss, Reward Accuracy, and Reward Margin during the training process are shown in Figure 1. It can be observed that the Training Loss decreased relatively steadily, while the Reward Accuracy and Reward Margin increased and then remained relatively stable.
(a) Training loss of DPO.
(b) Training reward accuracy of DPO.
(c) Training reward margin of DPO.
Figure 1: Training curves in DPO.
IV Evaluation
We conducted a comprehensive and rigorous evaluation of SynthCoder and other baseline models on both file-level and repository-level code completion tasks. Due to constraints in computational resources and time, we selected representative open-source models with Fill-in-the-Middle (FIM) capabilities that were released between 2024 and the first half of 2025. Models released earlier, such as CodeLlama-7b-hf  [^13] https://huggingface.co/codellama/CodeLlama-7b-hf from 2023, were excluded. Additionally, we excluded models not specifically fine-tuned for FIM tasks, such as MagicCoder  [^14] https://huggingface.co/ise-uiuc/Magicoder-S-DS-6.7B. All models were deployed on a single local H800 80G PCIe GPU, and inference was performed using VLLM version 0.9.1  [^15] https://github.com/vllm-project/vllm. The final set of models used for comparison is as follows.
  • Qwen2.5-Coder-7B  [^16] https://huggingface.co/Qwen/Qwen2.5-Coder-7B  A 7B parameter code language model developed by Alibaba, designed for code understanding and generation, serving as the base model in its series.
  • Qwen2.5-Coder-7B-Instruct  [^17] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct  The instruction-tuned version of the Qwen2.5-Coder-7B model.
  • deepseek-coder-6.7b-base  [^18] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base A 6.7B parameter base code model developed by DeepSeek, serving as the base model in its series.
  • deepseek-coder-6.7b-instruct  [^19] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct The instruction-tuned version of deepseek-coder-6.7b-base.
  • Qwen3-Coder-30B-A3B-Instruct  [^20] https://huggingface.co/Qwen/Qwen3-Coder-30B-A3B-Instruct An efficient code model from Alibaba’s Qwen3 series, which employs a Mixture-of-Experts (MoE) architecture and possesses excellent instruction-following capabilities.
  • Seed-Coder-8B-Base  [^21] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Base An 8B parameter base code model developed by the ByteDance Seed team. The model was trained using a model-driven data selection method, aiming to efficiently enhance its code generation abilities.
  • Seed-Coder-8B-Instruct  [^22] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Instruct The instruction-tuned version of Seed-Coder-8B-Base.
  • DeepSeek-Coder-V2-Lite-Base  [^23] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Base A lightweight version of DeepSeek-Coder-V2  [^24] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Base/, serving as a base model. It also utilizes a Mixture-of-Experts (MoE) architecture.
  • DeepSeek-Coder-V2-Lite-Instruct  [^25] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct The instruction-tuned version of DeepSeek-Coder-V2-Lite-Base.
  • StarCoder2-7B  [^26] https://huggingface.co/bigcode/starcoder2-7b A 7B parameter code model developed by the BigCode project. It was trained on The-Stack-V2, a large-scale dataset containing multiple programming languages, with a focus on code generation tasks.
  • aiXcoder-7b-v2  [^27] https://huggingface.co/aiXcoder/aixcoder-7b-v2 A 7B parameter code model designed to fully leverage long-context information, particularly excelling at cross-file code completion.
To ensure the rigor of our experimental results, we only included models in our comparison that were evaluated in the aforementioned file-level FIM code completion benchmarks and for which the official developers provided a specific FIM prompt format in their technical reports or open-source repository documentation.
  • The Qwen-Coder series (including Qwen2.5-Coder and Qwen3-Coder) models follow the format specified in their technical report [8];
  • The DeepSeek-Coder series models follow the example provided in the official repository  [^28] https://github.com/deepseek-ai/DeepSeek-Coder;
  • The Seed-Coder series models follow the official response in an issue on their official open-source repository  [^29] https://github.com/ByteDance-Seed/Seed-Coder/issues/4;
  • The StarCoder2 model follows the format described in its technical report [35];
  • The aiXcoder-7B-v2 model’s technical report [10] and open-source repository  [^30] https://github.com/aixcoder-plugin/aixcoder-7b-v2 only provide a prompt format for file-level completion and do not specify how to pass cross-file context. Therefore, we only evaluated this model in the file-level code completion tasks;
To conserve space in the data tables, we abbreviate Instruct to Inst and DeepSeek to DS. To validate the impact of different training stages and data compositions, we evaluated three versions of our model: SynthCoder-8B-SFT, SynthCoder-8B-DPO-wo-RS, and SynthCoder-8B-DPO-RS (the final version, referred to as SynthCoder-8B). We compared these against the base model, Seed-Coder-8B-Base. The differences between these model versions are detailed in Table I. To save space, all numerical values in tables are rounded to one decimal place. The percentage of increase or decrease refers to the change relative to SynthCoder-8B-DPO-RS.
TABLE I: Comparisons between SynthCoder and its varients.
Model Finetune Stage Alignment Stage Repetition Suppression
Seed-Coder-8B-Base NO NO NO
SynthCoder-8B-SFT YES NO NO
SynthCoder-8B-DPO-wo-RS YES YES NO
SynthCoder-8B-DPO-RS YES YES YES
IV-A File Level Completion
We select aiXcoder dataset as the evaluation dataset to verify our model’s capacity on file level code completetion.
We utilized the dataset  [^31] https://github.com/aixcoder-plugin/aiXcoder-7B/tree/main/aiXcoder_ex-tended_dataset provided by [11] to evaluate the model’s Fill-in-the-Middle (FIM) code completion performance on files from real-world open-source projects. The specific evaluation metrics used were Exact Match (EM) and Edit Similarity (ES). Our implementation is based on the reproduction project of Qwen Coder  [^32] https://github.com/QwenLM/Qwen3-Coder/tree/main/qwencoder-eval. We filtered out test cases where the total length of the context and the expected completion exceeded 8192 tokens (as tokenized by Qwen2.5-Coder-7B-Instruct) or where the generation length was greater than 64 tokens. During evaluation, the temperature was set to 0, and the maximum output length of the model was capped at 64 tokens.
As shown in Table II, SynthCoder-8B achieves an improvement of over 10% after finetuning, surpassing all baseline models, including those with a larger total parameter count such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B. Specifically, the fine-tuned SynthCoder-8B-SFT demonstrates an overall improvement of more than 10% compared to the base model, with a maximum gain of 22% on the JavaScript EM score. The aligned SynthCoder-8B-DPO-wo-RS shows further modest gains across all metrics, improving upon SynthCoder-8B-SFT by 4.2% in overall EM and 1.6% in overall ES. The final model with repetition suppression, SynthCoder-8B-DPO-RS, trades a slight decrease of up to 1.2% in C++ ES and on Python metrics for a 2.1% increase in overall EM when compared to SynthCoder-8B-DPO-wo-RS. Ultimately, this results in a total improvement of 17.8% in overall EM and 12.3% in overall ES relative to the base model.
TABLE II: Performance of models on aiXcoder dataset [11].
Model C++ JavaScript Java Python Overall
EM ES EM ES EM ES EM ES EM ES
aiXcoder-7B-v2 41.5 77.3 38.0 74.6 50.5 82.4 35.2 71.9 42.2 77.1
StarCoder2-7B 12.9 56.1 4.3 52.0 4.6 54.0 17.6 59.6 9.1 55.2
DS-Coder-6.7B-Inst 37.7 70.4 33.0 66.2 42.9 74.2 36.0 68.9 37.9 70.3
DS-Coder-V2-Lite-Inst 44.3 77.0 36.3 68.4 47.3 77.6 38.8 72.9 42.1 74.1
DS-Coder-6.7B-Base 45.4 76.8 40.4 73.5 52.6 81.9 42.8 75.4 46.0 77.4
DS-Coder-V2-Lite-Base 47.4 79.4 41.8 76.8 53.0 83.3 42.1 77.1 46.6 79.5
Qwen2.5-Coder-7B-Inst 38.8 70.6 34.4 67.1 49.0 77.4 33.0 66.2 39.8 71.0
Qwen3-Coder-30B-A3B 47.9 80.2 44.2 76.0 53.9 82.2 43.8 75.3 48.0 78.7
Qwen2.5-Coder-7B 49.3 79.9 44.1 77.6 53.6 82.8 45.0 78.3 48.5 80.0
Seed-Coder-8B-Inst 38.6 $\downarrow$28.2% 72.0 $\downarrow$12.8% 33.1 $\downarrow$37.4% 67.2 $\downarrow$17.2% 44.0 $\downarrow$28.4% 74.5 $\downarrow$15.3% 38.6 $\downarrow$16.9% 70.5 $\downarrow$11.2% 39.0 $\downarrow$27.6% 71.3 $\downarrow$14.3%
Seed-Coder-8B-Base 44.9 $\downarrow$10.2% 75.3 $\downarrow$7.8% 35.0 $\downarrow$30.0% 66.9 $\downarrow$17.8% 46.7 $\downarrow$20.9% 75.3 $\downarrow$14.1% 41.6 $\downarrow$8.4% 72.5 $\downarrow$8.3% 42.3 $\downarrow$17.8% 72.6 $\downarrow$12.3%
SynthCoder-8B-SFT 46.3 $\downarrow$6.7% 79.3 $\downarrow$2.4% 42.1 $\downarrow$8.0% 77.5 $\downarrow$1.6% 52.8 $\downarrow$6.9% 83.9 $\downarrow$2.4% 43.7 $\downarrow$3.2% 78.3 $\downarrow$0.1% 46.9 $\downarrow$6.3% 80.2 $\downarrow$1.7%
SynthCoder-8B-DPO-wo-RS 48.8 $\downarrow$1.4% 81.4 $\uparrow$0.3% 43.5 $\downarrow$4.7% 78.5 $\downarrow$0.4% 55.1 $\downarrow$2.4% 85.2 $\downarrow$0.8% 45.2 $\uparrow$0.2% 79.4 $\uparrow$1.2% 48.8 $\downarrow$2.1% 81.5 $\downarrow$0.1%
SynthCoder-8B-DPO-RS 49.5 81.2 45.5 78.8 56.5 85.9 45.1 78.5 49.8 81.5
IV-B Repository Level Code Completion
We selected ExecRepoBench, CrossCodeEval, and CoLT to evaluate the LLM’s capabilities in repository-level code completion.
IV-B1 CrossCodeEval with BM25 Retrieval Strategy
CrossCodeEval is a benchmark designed to evaluate the ability of LLMs to leverage cross-file context for code completion. Its data is sourced from open-source GitHub projects and covers four programming languages: Python, Java, TypeScript, and C#. The benchmark employs static analysis methods to construct test cases that are solvable only with the aid of cross-file context [14]. In this paper, we adopt the same evaluation metrics as the original paper – EM, ES, and ES_repoeval (abbreviated as ES_R) – and use the official cross-file context data provided by CrossCodeEval, which is retrieved using the BM25 text similarity algorithm. Each retrieved code snippet is limited to a maximum of 10 lines, and the top 5 most relevant snippets are selected. For each model, we utilize its specific prompt format for completion, with a input context length of 8192 tokens, a maximum output length of 50 tokens, and a temperature of 0.
As shown in Table III, SynthCoder-8B demonstrates a modest overall improvement compared to its base model and achieves the best performance among all models tested. This includes models with a larger parameter count, such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B. Specifically, the fine-tuned SynthCoder-8B-SFT shows an overall increase of 1.1% in EM, 0.6% in ES, and 0.7% in ES_R. While there was a slight decrease in the metrics for Java, all other languages showed improvement. The aligned SynthCoder-8B-DPO-wo-RS achieves gains of 0.6% to 0.7% overall, and also secures the highest scores on Java within the SynthCoder-8B series. The final version with repetition suppression, SynthCoder-8B-DPO-RS, obtains slight improvements of 1.5%, 1.5%, and 1.1% in overall EM, ES, and ES_R, respectively, when compared to the base model.
TABLE III: Performance of models on CrossCodeEval with BM25 retrieval strategy [14].
Model C# Java Python TypeScript Overall
EM ES ES_R EM ES ES_R EM ES ES_R EM ES ES_R EM ES ES_R
StarCoder2-7B 8.8 75.8 65.2 7.5 71.2 60.4 11.4 64.8 52.9 7.5 73.8 63.3 8.8 71.2 60.2
DS-Coder-6.7B-Inst 58.5 87.3 83.2 50.3 83.7 79.4 45.0 80.2 74.5 48.9 84.3 79.4 49.9 83.6 78.8
DS-Coder-V2-Lite-Inst 57.8 87.3 82.8 48.1 82.8 77.6 44.1 80.0 73.6 46.7 83.0 77.1 48.3 83.0 77.3
DS-Coder-6.7B-Base 61.8 88.2 84.7 51.8 84.2 80.1 46.9 80.7 75.2 50.2 84.2 79.4 51.7 84.0 79.3
DS-Coder-V2-Lite-Base 59.2 86.9 83.0 50.1 83.5 78.9 45.7 79.9 74.3 47.5 83.0 77.8 49.7 83.0 78.0
Qwen2.5-Coder-7B-Inst 46.7 83.4 78.0 46.6 81.3 76.5 41.8 78.2 72.1 44.3 81.5 76.1 44.6 80.9 75.4
Qwen3-Coder-30B-A3B 58.9 88.0 83.7 37.1 78.3 71.3 34.1 73.9 66.2 44.4 81.6 75.6 42.7 80.0 73.6
Qwen2.5-Coder-7B 61.5 87.8 84.2 50.1 83.7 79.3 45.4 79.8 74.2 47.9 83.7 78.6 50.1 83.4 78.6
Seed-Coder-8B-Inst 60.2 $\downarrow$8.9% 88.7 $\downarrow$2.8% 84.8 $\downarrow$3.3% 49.1 $\downarrow$5.0% 82.6 $\downarrow$3.0% 78.2 $\downarrow$2.8% 48.6 $\downarrow$1.9% 81.6 $\downarrow$1.5% 76.3 $\downarrow$1.1% 48.5 $\downarrow$6.5% 84.0 $\downarrow$2.8% 79.0 $\downarrow$2.7% 50.7 $\downarrow$5.5% 83.9 $\downarrow$2.5% 79.1 $\downarrow$2.4%
Seed-Coder-8B-Base 61.9 $\downarrow$5.9% 89.1 $\downarrow$2.3% 85.4 $\downarrow$2.5% 52.6 $\uparrow$1.9% 84.5 $\downarrow$0.8% 80.4 -0.0% 49.3 $\downarrow$0.3% 81.9 $\downarrow$1.2% 76.8 $\downarrow$0.4% 50.6 $\downarrow$1.9% 84.8 $\downarrow$1.8% 80.0 $\downarrow$1.4% 52.7 $\downarrow$1.5% 84.7 $\downarrow$1.5% 80.2 $\downarrow$1.1%
SynthCoder-8B-SFT 63.5 $\downarrow$3.3% 89.8 $\downarrow$1.4% 86.3 $\downarrow$1.4% 52.0 $\uparrow$0.8% 84.4 $\downarrow$0.9% 80.2 $\downarrow$0.2% 49.8 $\uparrow$0.6% 82.4 $\downarrow$0.6% 77.4 $\uparrow$0.3% 51.5 -0.0% 85.5 $\downarrow$1.0% 80.8 $\downarrow$0.4% 53.3 $\downarrow$0.4% 85.2 $\downarrow$0.9% 80.7 $\downarrow$0.4%
SynthCoder-8B-DPO-wo-RS 65.4 $\downarrow$0.2% 91.1 -0.0% 87.5 $\downarrow$0.1% 52.9 $\uparrow$2.5% 85.4 $\uparrow$0.3% 80.8 $\uparrow$0.5% 48.7 $\downarrow$1.5% 82.4 $\downarrow$0.6% 76.6 $\downarrow$0.8% 51.7 $\uparrow$0.3% 86.1 $\downarrow$0.3% 80.9 $\downarrow$0.2% 53.6 $\uparrow$0.2% 85.8 $\downarrow$0.2% 80.9 $\downarrow$0.2%
SynthCoder-8B-DPO-RS 65.5 91.1 87.6 51.6 85.1 80.4 49.5 82.9 77.2 51.6 86.3 81.1 53.5 86.0 81.0
IV-B2 CoLT with Dependency-based Strategy Only
The test set of CoLT-132K is a benchmark proposed in the aiXcoder-7B-v2 [10] to evaluate a model’s ability to utilize cross-file context for code completion. It encompasses Python, Java, C/C++, and Go with three completion scenarios: API calls, single-line completions, and span completions, comprising a total of 12,000 samples.
In this work, we use Exact Match (EM) and Edit Similarity (ES) as the evaluation metrics. To specifically assess the model’s ability to leverage dependency-based cross-file context, we exclusively retain the context retrieved through dependency call relationships for each test case. For each model, we employ its corresponding prompt format for completion, with standardized total input context length of 4096 tokens, maximum output length of 200 tokens, and temperature of 0.
As demonstrated in Table IV, SynthCoder-8B shows improvement over its base model in all four programming languages and achieves the best performance among all tested models. This includes outperforming models with larger parameter counts, such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B. The fine-tuned SynthCoder-8B-SFT improves the overall EM and ES by 4.1% and 4.2%, respectively, with gains across all individual metrics, including a maximum improvement of 7.4% on the Go EM score. The aligned SynthCoder-8B-DPO-wo-RS further enhances performance over SynthCoder-8B-SFT on all metrics, with an additional increase of 1.5% in overall EM and 1.4% in overall ES. The final version with repetition suppression, SynthCoder-8B-DPO-RS, achieves the optimal results across all metrics, ultimately boasting a 7.0% improvement in overall EM and a 7.5% improvement in overall ES compared to the base model. It is worth noting that for all metrics on this benchmark, each subsequent stage of our process – fine-tuning, alignment, and repetition suppression – resulted in a consistent improvement in the performance of the SynthCoder-8B model.
TABLE IV: Performance of models on CoLT with dependency-based strategy only [10].
Model C++ Go Java Python Overall
EM ES EM ES EM ES EM ES EM ES
StarCoder2-7B 32.2 59.8 33.5 62.7 44.5 66.0 29.2 59.6 34.8 62.0
DS-Coder-6.7B-Inst 37.7 65.9 47.0 75.4 53.8 75.9 37.5 67.6 44.1 71.2
DS-Coder-V2-Lite-Inst 40.4 69.2 49.2 77.2 56.4 79.7 40.3 73.0 46.6 74.8
DS-Coder-6.7B-Base 41.0 69.3 51.6 79.4 57.6 79.9 40.9 71.8 47.9 75.1
DS-Coder-V2-Lite-Base 41.7 70.8 51.9 79.1 58.0 81.6 42.9 73.9 48.6 76.4
Qwen2.5-Coder-7B-Inst 42.3 71.0 47.4 76.3 54.8 78.5 39.2 71.0 45.9 74.2
Qwen3-Coder-30B-A3B 47.7 73.5 52.3 78.9 58.1 78.4 46.1 74.5 51.1 76.3
Qwen2.5-Coder-7B 35.8 58.5 47.6 74.7 47.7 65.1 38.5 66.5 42.4 66.2
Seed-Coder-8B-Inst 42.8 $\downarrow$16.0% 71.7 $\downarrow$12.1% 48.5 $\downarrow$15.2% 76.2 $\downarrow$9.9% 57.7 $\downarrow$8.2% 81.1 $\downarrow$6.3% 46.0 $\downarrow$12.5% 75.1 $\downarrow$8.3% 48.8 $\downarrow$12.7% 76.0 $\downarrow$9.1%
Seed-Coder-8B-Base 46.4 $\downarrow$7.0% 74.2 $\downarrow$8.3% 51.0 $\downarrow$9.7% 76.6 $\downarrow$9.3% 59.6 $\downarrow$4.6% 81.6 $\downarrow$5.7% 48.4 $\downarrow$6.9% 76.3 $\downarrow$6.6% 51.4 $\downarrow$7.0% 77.2 $\downarrow$7.5%
SynthCoder-8B-SFT 47.5 $\downarrow$4.5% 76.6 $\downarrow$4.9% 54.7 $\downarrow$2.3% 81.3 $\downarrow$3.1% 60.8 $\downarrow$2.6% 83.8 $\downarrow$2.9% 50.5 $\downarrow$2.4% 79.4 $\downarrow$2.5% 53.4 $\downarrow$2.9% 80.2 $\downarrow$3.3%
SynthCoder-8B-DPO-wo-RS 48.9 $\downarrow$1.6% 78.6 $\downarrow$2.2% 55.1 $\downarrow$1.6% 82.0 $\downarrow$2.1% 62.1 $\downarrow$0.5% 85.3 $\downarrow$1.1% 50.6 $\downarrow$2.2% 79.6 $\downarrow$2.2% 54.2 $\downarrow$1.4% 81.4 $\downarrow$1.9%
SynthCoder-8B-DPO-RS 49.7 80.3 55.9 83.8 62.4 86.2 51.7 81.4 54.9 82.9
IV-B3 ExecRepoBench
The ExecRepoBench benchmark  [^33] https://huggingface.co/datasets/CSJianYang/ExecRepoBench, proposed by the Alibaba team, focuses on evaluating the code completion capabilities of LLMs in real-world software development environments. The dataset is sourced from 50 actively maintained Python repositories on GitHub, comprising a total of 1,200 test cases. For test case construction, the benchmark introduces a multi-level syntax masking method based on Abstract Syntax Trees (AST). It utilizes Tree-sitter to classify syntax nodes into three categories – functions, statements, and expressions – to create test instances. Additionally, it employs non-syntactic construction methods, including snippet, single-line, and multi-line truncation. The benchmark provides not only the intra-file prefix and suffix in a Fill-in-the-Middle (FIM) format but also the relevant cross-file context from the same repository in the form of complete files.
In our reproduction of the ExecRepoEval evaluation, we implemented a rule-based fix to correct indentation errors in the code generated by the models. We use Pass@1 and Edit Similarity (ES) to measure model performance. For each model, we employed its corresponding prompt format for completion, with standardized limits of 4096 tokens for the intra-file context, 4096 tokens for the cross-file context, and a maximum output length of 160 tokens. The temperature was set to 0.
As shown in Table V, SynthCoder-8B achieves a 24.5% improvement in Pass@1 and a 35.3% improvement in ES compared to its base model. Its performance is also significantly higher than that of other 7B and 14B models, including Qwen2.5-Coder-14B-Inst. We did not include Qwen2.5-Coder-Instruct-C, which was proposed in the original ExecRepoBench paper, as its weights have not been publicly released. The fine-tuned SynthCoder-8B-SFT improves upon the base model by 8.4% in Pass@1 and 8.6% in ES. The aligned SynthCoder-8B-DPO-wo-RS further boosts performance by 15.5% in Pass@1 and 25.1% in ES relative to SynthCoder-8B-SFT. The final version with repetition suppression, SynthCoder-8B-DPO-RS, adds an additional 0.6% to Pass@1 and 1.6% to ES, culminating in a total improvement of 24.5% and 35.3% over the original base model.
TABLE V: Performance of models on ExecRepoEval [12].
Model ES Pass@1
StarCoder2-7B 18.9 21.8
DS-Coder-6.7B-Inst 23.3 32.4
DS-Coder-V2-Lite-Inst 46.4 41.1
DS-Coder-6.7B-Base 39.4 42.1
DS-Coder-V2-Lite-Base 42.5 39.6
Qwen2.5-Coder-7B-Inst 36.2 39.3
Qwen3-Coder-30B-A3B 35.2 41.5
Qwen2.5-Coder-7B 18.2 19.4
Seed-Coder-8B-Inst 32.8 $\downarrow$89.3% 38.1 $\downarrow$36.0%
Seed-Coder-8B-Base 45.9 $\downarrow$35.3% 41.6 $\downarrow$24.5%
SynthCoder-8B-SFT 49.0 $\downarrow$26.7% 44.6 $\downarrow$16.1%
SynthCoder-8B-DPO-wo-RS 61.1 $\downarrow$1.6% 51.5 $\downarrow$0.6%
SynthCoder-8B-DPO-RS 62.1 51.8
IV-C Errors of Wrongly Repeating Suffix or Prefix
When applying base models to practical code completion plugins, we observed that models frequently tend to erroneously repeat the preceding (prefix) or succeeding (suffix) context. This behavior can frustrate users and lead them to question the capabilities of both the model and the plugin. To analyze this issue systematically, we focus on a specific scenario: cases where the first non-empty line of the generated content does not exactly match the expected completion, but is identical to the last line of the prefix or the first line of the suffix after whitespace is removed. Our models significantly reduce the probability of such anomalous repetitions. This definition covers instances where the model repeats a context line and then proceeds to generate additional code. Although APIs like OpenAI’s provide various repetition penalty parameters, their effectiveness is limited in the context of Fill-in-the-Middle (FIM) tasks. This is due to the substantially longer input lengths compared to output lengths and the model’s extremely low perplexity when it erroneously repeats the context. To suppress this problem thoroughly, approaches like model-side training or output interception are required.
TABLE VI: Suffix repetition ratio by model and dataset.
Model SC* AX* ERE* CoLT
StarCoder2-7B 0.33% 1.38% 30.58% 1.87%
DS-Coder-6.7B-Inst 0.13% 1.62% 4.90% 1.98%
DS-Coder-V2-Lite-Inst 0.15% 0.64% 2.15% 2.19%
DS-Coder-6.7B-Base 0.13% 0.77% 4.73% 2.10%
DS-Coder-V2-Lite-Base 0.15% 1.00% 4.64% 2.11%
Qwen2.5-Coder-7B-Inst 0.13% 3.23% 1.03% 2.58%
Qwen3-Coder-Flash 0.10% 1.16% 0.26% 1.62%
Qwen2.5-Coder-7B 0.23% 0.69% 0.09% 1.51%
Seed-Coder-8B-Inst 0.31% $\uparrow$210% 2.74% $\uparrow$621% 0.43% $\uparrow$377% 2.54% $\uparrow$418%
Seed-Coder-8B-Base 0.17% $\uparrow$70% 1.67% $\uparrow$339% 2.23% $\uparrow$2377% 1.77% $\uparrow$261%
SynthCoder-SFT 0.10% -0% 0.60% $\uparrow$57% 2.58% $\uparrow$2766% 1.42% $\uparrow$189%
SynthCoder-DPO-wo-RS 0.11% $\uparrow$10% 0.52% $\uparrow$36% 0.60% $\uparrow$566% 0.50% $\uparrow$2%
SynthCoder-DPO-RS 0.10% 0.38% 0.09% 0.49%
    SC: SantaCoder; AX: aiXcoder; ERE: ExecRepoEval.
TABLE VII: Prefix repetition ratio by model and dataset.
Model SC* AX* ERE* CoLT
StarCoder2-7B 0.00% 0.49% 0.17% 0.04%
DS-Coder-6.7B-Inst 0.00% 0.43% 0.17% 0.03%
DS-Coder-V2-Lite-Inst 0.00% 0.41% 0.60% 0.08%
DS-Coder-6.7B-Base 0.00% 0.57% 0.34% 0.06%
DS-Coder-V2-Lite-Base 0.00% 0.52% 0.95% 0.08%
Qwen2.5-Coder-7B-Inst 0.00% 0.17% 0.43% 0.04%
Qwen3-Coder-Flash 0.00% 0.15% 0.09% 0.04%
Qwen2.5-Coder-7B 0.00% 0.18% 0.43% 0.04%
Seed-Coder-8B-Inst 0.00% -0% 0.33% $\uparrow$560% 0.09% -0% 0.08% $\uparrow$300%
Seed-Coder-8B-Base 0.00% -0% 0.56% $\uparrow$1020% 0.09% -0% 0.08% $\uparrow$300%
SynthCoder-SFT 0.00% -0% 0.15% $\uparrow$200% 0.34% $\uparrow$278% 0.16% $\uparrow$700%
SynthCoder-DPO-wo-RS 0.00% -0% 0.10% $\uparrow$100% 0.17% $\uparrow$89% 0.08% $\uparrow$300%
SynthCoder-DPO-RS 0.00% 0.05% 0.09% 0.02%
    SC: SantaCoder; AX: aiXcoder; ERE: ExecRepoEval.
We analyzed the frequency of erroneous suffix and prefix repetitions in the generated outputs across the aforementioned datasets, with the results presented in Table VI and Table VII, respectively. The CrossCodeEval dataset is omitted from these tables because its ground-truth completions do not contain full lines, meaning the specific type of whole-line repetition we define does not occur. We observed that for models using either the Prefix-Suffix-Middle (PSM) FIM format (e.g., the Qwen-Coder and DeepSeek series) or the Suffix-Prefix-Middle (SPM) FIM format (e.g., the Seed-Coder series), the frequency of erroneously repeating the suffix is generally higher than that of repeating the prefix. The StarCoder2-7B model exhibits a particularly high repetition rate, reaching a suffix repetition rate of 30.58% on the ExecRepoEval benchmark. While the prefix repetition rate of DeepSeek-Coder is generally lower than that of the Qwen-Coder series, the opposite is true for their suffix repetition rates. The Seed-Coder series of models demonstrates lower repetition rates compared to the other model families.
We also noted that comparing the model before and after fine-tuning (i.e., Seed-Coder-8B-Base vs. SynthCoder-SFT), the repetition rates exhibit fluctuations of varying magnitudes and directions across different test sets. Interestingly, even without the inclusion of specialized repetition-suppression data, the context repetition rate of SynthCoder-DPO-wo-RS is lower than that of SynthCoder-SFT. This is because outputs containing repetitions that are generated and rejected during the sampling process are subsequently used as the dispreferred examples in the preference pairs for DPO training. After incorporating an additional 11% of repetition-suppression data, SynthCoder-DPO-RS demonstrated a significantly greater reduction in this behavior. It achieved a prefix repetition rate below 0.5% and a suffix repetition rate below 0.1% across all datasets, marking the best performance among the evaluated models.
V Discussion
In this section, we discuss some implications of our study.
V-A Gap Between Benchmarks and Application
Section IV-C vividly illustrates the persistent gap between a model’s strong performance on benchmark metrics and its practical utility in a production environment. Truncation strategies employed during evaluation may not align with real-world scenarios, and conversely, benchmark scores often fail to capture the nuances of user experience and actual code effectiveness. The issue of erroneous context repetition is a prime example. While a straightforward solution is to filter such suggestions at the client-side (i.e., in the code completion tool), this approach does not prevent the model from expending computational resources on generating the flawed output. A more fundamental solution lies in refining the model to produce more desirable content directly – either by avoiding such outputs altogether or by optimizing the decoding process. Furthermore, models can generate content entirely irrelevant to code completion. For instance, Qwen3-Coder-30B-A3B-Instruct often misinterprets its task as generating a Markdown code block, including not only the ‘‘‘ delimiters but also explanatory text. Another concern is the inadvertent disclosure of private information, such as organization or user names, which can be problematic for users. Addressing these challenges requires a concerted effort, spanning from the curation of training data on the model side to the implementation of output filtering mechanisms on the tool side.
We also observed significant variations in output length among different models given the same input. Ideally, generated code should be concise and to the point, but this is often not the case. Models without fine-tuning are prone to generating excessively long outputs, either by endlessly repeating permutations of the surrounding code or by arbitrarily adding extensive, non-essential docstrings. This can lead to the model reaching its maximum generation length, resulting in unusable code. This behavior is particularly pronounced in the Qwen-Coder series of models and is frequently observed in the ExecRepoEval benchmark. For developers, such unbounded generation not only needlessly consumes computational resources on the model side but also wastes their time and mental energy waiting for and evaluating the results. This underscores the importance of constructing training corpora based on syntax tree nodes and user behavior patterns, which enables the model to generate more targeted and appropriately sized completions. These crucial capabilities are not adequately captured by current evaluation benchmarks.
In summary, enhancing the performance of code completion necessitates a collaborative effort between LLM and client-side.
V-B About the Context Ranges
Is a longer context always better in code completion? In the work on HCP, the team demonstrated through experiments that removing global context information from the cross-file context improves the model’s completion performance compared to no pruning [45]. Similarly, aiXcoder [10] found that in repository-level code completion, LLMs may overlook relevant APIs and similar code snippets within the input context. These findings suggest that a longer context is not axiomatically superior; a complete project-level codebase may contain distracting information, and models may not be able to directly identify and leverage all the effective information present. Recent work has begun to integrate static analysis techniques not just into data construction or online inference, but also into the training process. For example, the STALL+ framework [33] injects static analysis at three critical stages: 1) during prompting, to extract file-level API signatures and a set of valid symbol-level tokens; 2) during decoding, to constrain generation to legal API calls by setting the probabilities of illegal tokens to zero; and 3) during post-processing, to validate the output with a compiler, achieving significant improvements on identifier matching metrics.
In conclusion, for project-level code completion, filtering relevant information from the cross-file context before inputting it to an LLM yields better results than simply extending the number or total length of context snippets. LLMs should also learn to adaptively select the appropriate context based on the specific inference scenario. In this work, we extract cross-file context from the dual perspectives of similarity and dependency. By retaining only highly similar code fragments and essential information such as signature definitions, we significantly increase the density of pertinent information in the input. Moreover, we maintain consistency between the data filtering methods used during training and those applied during inference, which enhances the project-level code completion capabilities of SynthCoder in practical application scenarios.
V-C Next Edit Suggestions
Next Edit Prediction infers a developer’s intent by analyzing the code context and historical edit sequences. It can predict the content and location of the next edit without requiring natural language input, thereby preventing interruptions to the workflow [46]. For instance, during repetitive tasks such as API updates or identifier renaming, it can proactively suggest coherent subsequent edits based on the user’s initial modification. This feature has already been implemented in GitHub Copilot  [^34] https://code.visualstudio.com/blogs/2025/02/12/next-edit-suggestions, where it works in synergy with traditional code completion capabilities.
The core of the edit prediction task is to modify existing code: whether to fix bugs or iterate on features, rather than to rewrite it entirely. This area has been a subject of extensive academic research. For example, CoditT5 [47] introduced a pre-trained model specifically designed for code editing. Modit [48] developed a multi-modal editing engine that integrates edit location, code context, and commit information. More recent approaches like jLED [49] employ a joint training methodology, enabling the model to simultaneously learn where to edit and how to edit, providing a more practical solution for real-world editing scenarios. Collectively, these research efforts have advanced the ability of models to follow precise instructions and execute complex code modifications.
In contrast to the fine-grained task of performing insertions, deletions, and modifications at the cursor, many current methods for evaluating the coding capabilities of large models focus on code generation, which does not fully capture the nuances of edit prediction. For example, mainstream function-level benchmarks, including HumanEval [50] and MBPP [51], require models to generate complete code blocks based on function signatures, docstrings, or functional descriptions. The limitation of such benchmarks is that they measure a model’s ability to comprehend high-level instructions and generate code from scratch, rather than its capacity for making precise, intelligent, and incremental modifications within the complex context of existing code. Meanwhile, existing code completion datasets primarily focus on token-level completion at the cursor based on the surrounding code context; they do not incorporate edit history, making it difficult to assess a model’s ability to reason by analogy from past edits. Looking ahead, models such as zeta  [^35] https://huggingface.co/zed-industries/zeta have already successfully integrated FIM-based code completion with edit prediction, achieving consistently strong performance across both capabilities. This integration will enable models to become more intelligent and better aligned with developer workflows, driving a paradigm shift in code completion from a reactive response to a proactive prediction model that incorporates edit suggestions.
VI Threats to Validity
In this section, we discuss some factors that may potentially bring risks to the findings of the study.
Limited models
Actually, we established SynthCoder based on only Seed-Coder-8B-Base as it is the current state-of-the-art (SOTA) LLM in code completion so far. However, SynthCoder is designed as a model-agnostic approach, therefore any LLM with capability of conducting FIM code completion tasks can be post-trained to create variations of SynthCoder. In this sense, it raises the necessity to replicate our approaches using other LLM series.
Limited datasets
The evaluation of SynthCoder is based on a limited number of datasets and benchmark methods, which may bring in certain validity risks. In particular, the test sets may not fully cover the diversity of real-world application scenarios, leading to potential gaps between benchmark results and practical performance. However, we have alreadly selected industry-recognized open-source benchmarks (ExecRepoEval and CrossCodeEval) and a custom test set constructed from third-party open-source datasets. Meanwhile, we evaluated the capabilities of SynthCoder and other baseline models in both single-file code completion, cross-file code completion based on dependency graph retrieval, and cross-file code completion based on BM25 retrieval so as to establish a comprehensive understanding of the performance of SynthCoderṪhe results strongly demonstrate the superiority of our method and model. Throughout the evaluation process, we utilized the official testing frameworks and parameters provided by the benchmarks or other well-known open-source models wherever possible to ensure the validity and generalizability of our tests.
Data leakage
Data leakage seems to be an inevitable risk. It may lead to inflated evaluation results that are not representative of real-world performance. Nevertheless, we have taken corresponding measures to alleviate this risk. To detect potential data leakage, we recorded the source repository URLs and intra-repository file paths for all synthetic data derived from The-Stack-V2 that was used for training. We then compared this information against the repository and file paths recorded in the aiXcoder, ExecRepoEval, CrossCodeEval, and CoLT test sets. By constructing and comparing a repository-file tuple structure, we controlled the possibility of data contamination resulting from overlapping files between the training and test sets.
VII Conclusion
By integrating multiple best practices, we have fine-tuned the Seed-Coder-8B-Base to achieve state-of-the-art (SOTA) performance on the Fill-in-the-Middle code completion task for models with fewer than 10B active parameters. Furthermore, we discuss optimization directions for deploying such models in production environments that extend beyond conventional evaluation metrics.

### v2

@@
 As the Go language was not included in this subset, we used the synthetic_qa subset for extraction instead.
-Since functions in this dataset may lack docstrings, we used the DeepSeek-V3-0324  [^8] https://huggingface.co/deepseek-ai/DeepSeek-V3-0324 model to synthesize corresponding docstrings for the extracted functions.
+Since functions in this dataset may lack docstrings, we used the DeepSeek-V3-0324888https://huggingface.co/deepseek-ai/DeepSeek-V3-0324 model to synthesize corresponding docstrings for the extracted functions written in Go.
 III-A3 Filtering Based on Data Quality
@@
 For a sequence $x=(x_{1},x_{2},...,x_{N})$ of length $N$, where the model’s predicted conditional probability is $P(x_{t}|x_{<t})$, the perplexity $\text{PPL}(x)$ is as follows:
-$\text{PPL}(x)=\exp(-\frac{1}{N}\sum_{t=1}^{N}\log P(x_{t}|x_{<t}))$
-In this work, we compute the perplexity of the code completion and code infilling data on the LLM to be trained.
-The distribution of the perplexity values generally approximates a log-normal distribution.
-For the post-training process, data with very low perplexity represents knowledge the LLM has already mastered and can output correctly without further training.
-Conversely, data with very high perplexity deviates significantly from the distribution of the LLM’s base data, and this subset tends to have a higher proportion of erroneous data.
-We remove a certain percentage of the data with the lowest and highest perplexity to improve the overall quality of the training set [37, 38].
-For the infilling data, we ultimately remove data outside of two standard deviations (2 $\sigma$) of the fitted log-normal distribution.
-For the completion data, we discard the lowest and highest 5% of the data.
+$\text{PPL}(x)=\exp\Big{(}-\frac{1}{N}\sum_{t=1}^{N}\log P(x_{t}|x_{<t})\Big{)}$
+[…]
+Figure 1: Composition of the dataset by generation strategy.
 III-B Data Construction
@@
 (2) For building the completion context, we introduce critical cross-file information using code similarity (BM25) and code dependency graphs (Call Graphs).
-(3) For optimizing data quality, we employ rejection sampling and a multi-stage filtering process to generate high-quality preference pairs for alignment training, aiming to enhance the model’s FIM capabilities
-and suppress context repetition.
+(3) For optimizing data quality, we employ rejection sampling and a multi-stage filtering process to generate high-quality preference pairs for alignment training, aiming to enhance the model’s FIM capabilities and suppress context repetition.
 Through this multi-dimensional construction strategy, we aim to generate high-quality training data that more closely resembles real-world application scenarios.
 III-B1 Inner File Context
-Today, AI-driven coding tools like GitHub Copilot  [^10] https://github.com/features/copilot and Cursor  [^11] https://www.cursor.com have been widely adopted by the developer community.
-Because developers have certain inherent habits in their thinking and coding processes – habits which are in turn influenced and reshaped by these AI tools – the locations where they expect to trigger code completion tend to follow specific patterns.
-Based on this core idea, our work strives to construct code completion data that aligns as closely as possible with developer habits.
-We have designed multiple strategies for this purpose, which can be broadly categorized into two main types: a strategy based on Abstract Syntax Tree (AST) nodes (60% of the data) and a heuristic strategy based on user habits (40%).
+The widespread adoption of AI-driven coding tools has profoundly influenced the developer community.
+These tools interact with and shape developers’ inherent coding habits, leading to predictable patterns in the locations where they trigger code completion.
+Motivated by this observation, our work aims to construct a code completion dataset that closely mimics these real-world developer behaviors.
+To this end, we have designed multiple data generation strategies, which, as illustrated in Figure 1, are grouped into three primary categories: Strategies Based on AST Nodes (blue, abbreviated as AST), which constitute the majority at $66.89\%$; Strategies Based on User Behavior Patterns (orange, abbreviated as User), accounting for $22.56\%$; and Strategies Based on Random Single or Multiple Lines (green, abbreviated as Rand), which make up the remaining.
 Strategies Based on AST Nodes
@@
 Based on a variety of node extraction strategies, we select target nodes as the code snippets to be completed and extract their corresponding prefix and suffix contexts.
-Compared to previous related work [39, 11, 12], we categorize syntactically or functionally similar nodes together, thereby creating 13 more fine-grained data synthesis strategies.
-These are based on: classes and interfaces; methods; formal and actual parameters; imported modules; expressions; assignment operations and variable declarations; comments; conditional statements; loop statements; return statements; other common statements; annotations or decorators; and core concurrent programming statements (only in Go).
+To create fine-grained data, we identified 13 distinct strategies based on syntactically or functionally similar AST nodes, as detailed in the blue sections of Figure 1.
+These strategies vary in prevalence.
+The most significant contributions come from targeting locations based on Expressions ($10.66\%$), Other Common Statement Blocks ($9.90\%$), and Methods ($8.90\%$).
+Other key strategies address fundamental constructs such as assignment operations and variable declarations ($6.99\%$), parameters and arguments ($6.53\%$), and conditional statements ($5.43\%$).
+The remaining strategies cover more specific contexts, including annotations or decorators ($0.46\%$) and concurrent statements in Go ($0.18\%$), ensuring comprehensive coverage of syntactic structures.
 Strategies Based on User Behavior Patterns
@@
 Specifically, this includes the following three types of data.
-  • Code Completion Following a Comment: When a model cannot directly understand a developer’s intent, the developer may manually write a comment describing the functionality of the code block to be written, then trigger code completion on the next line.
-To simulate this, we use Tree-sitter to detect full-line comment nodes in code files and treat the first code block immediately following the comment as the completion target.
-  • Code Completion Within Parentheses: Paired parentheses often enclose relatively independent code snippets.
+  • Incomplete Single Lines (14.78%): When a developer pauses typing and there is no code to the right of the cursor, code completion is often triggered to finish the current line.
+We designed two strategies for this scenario.
+    • Random Intra-Line Completion: The completion is triggered at a randomly selected position within a line of code. The subsequent text on that line serves as the completion target.
+    • Syntax-Aware Triggered Completion: The completion is triggered immediately following a language-specific, syntactically significant token.
+We define a curated set of such tokens for each programming language, encompassing critical keywords (e.g., for control flow, definitions) and operators (e.g., for assignment, member access).
+The remainder of the line after the token is designated as the completion target.
+  • Code Fragments Within Parentheses (4.86%): Paired parentheses often enclose relatively independent code snippets.
 In code editors, typing an opening parenthesis automatically generates the closing one and places the cursor in between.
 We sample data to replicate this common scenario.
-  • Completion of Incomplete Single Lines: When a developer pauses typing and there is no code to the right of the cursor, code completion is often triggered to finish the current line.
-We designed two strategies for this scenario.
-    • Completion based on a random position: we randomly select a point that is not at the beginning or end of a line and designate the rest of the line as the completion target.
-    • Completion based on special symbols: we compile a list of specific keywords and special characters for each language, search for these tokens during construction, and use the remainder of the line as the target.
+  • Code Fragments After Comments (2.92%): When a model cannot directly understand a developer’s intent, the developer may manually write a comment describing the functionality of the code block to be written, then trigger code completion on the next line.
+To simulate this, we use Tree-sitter to detect full-line comment nodes in code files and treat the first code block immediately following the comment as the completion target.
+Strategies Based on Random Single or Multiple Lines
 To enhance the LLM’s generalization capabilities, we also randomly select single lines and consecutive multi-line blocks of code as completion targets, which accounts for approximately 10% of the total data volume.
@@
     We select the top $k\%$ (k=30) most complex fragments to serve as FIM and completion training samples, respectively, forming the final training dataset.
-Training Parameters and Cost
+Training Details
 In the curriculum learning fine-tuning stage, we trained the model for one epoch using the AdamW optimizer, with a learning rate set to 3e-6 combined with a 50-step warmup.
@@
 Strategies such as DeepSpeed ZeRO-1, Flash Attention, and gradient checkpointing activation re-computation were used to improve training efficiency.
+Finally, the process accounted for 800 steps and finished in 20 hours.
 III-C2 Alignment Stage
@@
 $\quad\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\Bigg{]}$
-Training Parameters and Cost
+Training Details
 In the preference alignment stage using the DPO algorithm, the DPO parameter $\beta$ was set to 0.9, with an equivalent batch size of 64.
@@
 The changes in Loss, Reward Accuracy, and Reward Margin during the training process are shown in Figure 2.
-It can be observed that the Training Loss decreased relatively steadily, while the Reward Accuracy and Reward Margin increased and then remained relatively stable.
+The alignment process lasted for 400 steps within 24 hours, and it can be observed that the Training Loss decreased relatively steadily, while the Reward Accuracy and Reward Margin increased and then remained relatively stable.
 (a) Training loss of DPO.
@@
 IV Evaluation
-We conducted a comprehensive and rigorous evaluation of SynthCoder and other baseline models on both file-level and repository-level code completion tasks.
-Due to constraints in computational resources and time, we selected representative open-source models with Fill-in-the-Middle (FIM) capabilities that were released between 2024 and the first half of 2025.
+We conducted a comprehensive and rigorous evaluation of SynthCoder and baseline models on both file-level and repository-level code completion tasks.
+Due to constraints in computational resources, we selected representative open-source models with Fill-in-the-Middle (FIM) capabilities that were released between 2024 and the first half of 2025.
 Models released earlier, such as CodeLlama-7b-hf [^11] https://huggingface.co/codellama/CodeLlama-7b-hf from 2023, were excluded.
 Additionally, we excluded models not specifically fine-tuned for FIM tasks, such as MagicCoder [^12] https://huggingface.co/ise-uiuc/Magicoder-S-DS-6.7B.
-All models were deployed on a single local H800 80G PCIe GPU, and inference was performed using VLLM version 0.9.1  [^15] https://github.com/vllm-project/vllm.
+All models were deployed on a single local H800 80G PCIe GPU, and inference was performed using VLLM version 0.9.1131313https://github.com/vllm-project/vllm.
 The final set of models used for comparison is as follows.
-  • Qwen2.5-Coder-7B  [^16] https://huggingface.co/Qwen/Qwen2.5-Coder-7B  A 7B parameter code language model developed by Alibaba, designed for code understanding and generation, serving as the base model in its series.
-  • Qwen2.5-Coder-7B-Instruct  [^17] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct  The instruction-tuned version of the Qwen2.5-Coder-7B model.
-  • deepseek-coder-6.7b-base  [^18] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base A 6.7B parameter base code model developed by DeepSeek, serving as the base model in its series.
-  • deepseek-coder-6.7b-instruct  [^19] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct The instruction-tuned version of deepseek-coder-6.7b-base.
-  • Qwen3-Coder-30B-A3B-Instruct  [^20] https://huggingface.co/Qwen/Qwen3-Coder-30B-A3B-Instruct An efficient code model from Alibaba’s Qwen3 series, which employs a Mixture-of-Experts (MoE) architecture and possesses excellent instruction-following capabilities.
-  • Seed-Coder-8B-Base  [^21] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Base An 8B parameter base code model developed by the ByteDance Seed team.
-The model was trained using a model-driven data selection method, aiming to efficiently enhance its code generation abilities.
-  • Seed-Coder-8B-Instruct  [^22] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Instruct The instruction-tuned version of Seed-Coder-8B-Base.
-  • DeepSeek-Coder-V2-Lite-Base  [^23] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Base A lightweight version of DeepSeek-Coder-V2  [^24] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Base/, serving as a base model.
-It also utilizes a Mixture-of-Experts (MoE) architecture.
-  • DeepSeek-Coder-V2-Lite-Instruct  [^25] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct The instruction-tuned version of DeepSeek-Coder-V2-Lite-Base.
-  • StarCoder2-7B  [^26] https://huggingface.co/bigcode/starcoder2-7b A 7B parameter code model developed by the BigCode project.
-It was trained on The-Stack-V2, a large-scale dataset containing multiple programming languages, with a focus on code generation tasks.
-  • aiXcoder-7b-v2  [^27] https://huggingface.co/aiXcoder/aixcoder-7b-v2 A 7B parameter code model designed to fully leverage long-context information, particularly excelling at cross-file code completion.
+  • aiXcoder-7b-v2141414https://huggingface.co/aiXcoder/aixcoder-7b-v2
+    A 7B parameter code LLM designed to fully leverage long-context information, particularly excelling at cross-file code completion.
+  • StarCoder2-7B [^15] https://huggingface.co/bigcode/starcoder2-7b
+    A 7B parameter code LLM developed by the BigCode project. It was trained on The-Stack-V2, a large-scale dataset containing multiple programming languages.
+  • DeepSeek-Coder-6.7B (-base [^16] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base, -instruct [^17] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct)
+    A pair of 6.7B code LLMs developed by DeepSeek. The former is the base model, while deepseek-coder-6.7b-instruct is the instruction-tuned version of deepseek-coder-6.7b-base.
+  • Qwen2.5-Coder-7B [^18] https://huggingface.co/Qwen/Qwen2.5-Coder-7B (-Instruct [^19] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct)
+    A pair of 7B code LLMs developed by Alibaba, designed for code understanding and generation.
+Qwen2.5-Coder-7B is the base model, while Qwen2.5-Coder-7B-Instruct is the instruction-tuned version of Qwen2.5-Coder-7B.
+  • Seed-Coder-8B (-Base [^20] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Base, -Instruct [^21] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Instruct)
+    A pair of 8B parameter code LLMs developed by the ByteDance.
+The model was trained using a model-driven data selection method, aiming to efficiently enhance its code generation abilities, while Seed-Coder-8B-Instruct is the instruction-tuned version of Seed-Coder-8B-Base.
+  • DeepSeek-Coder-V2-Lite (-Base [^22] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Base, -Instruct [^23] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct)
+    A pair of lightweight code LLMs in DeepSeek-Coder-V2 model family [^24] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Base.
+Both of them utilize a Mixture-of-Experts (MoE) architecture, while DeepSeek-Coder-V2-Lite-Instruct is the instruction-tuned version of DeepSeek-Coder-V2-Lite-Base.
+  • Qwen3-Coder-30B-A3B-Instruct [^25] https://huggingface.co/Qwen/Qwen3-Coder-30B-A3B-Instruct
+    An efficient instruction-tuned code model from Alibaba’s Qwen3 series, which also employs a Mixture-of-Experts (MoE) architecture and possesses excellent instruction-following capabilities.
 To ensure the rigor of our experimental results, we only included models in our comparison that were evaluated in the aforementioned file-level FIM code completion benchmarks and for which the official developers provided a specific FIM prompt format in their technical reports or open-source repository documentation.
@@
   • The aiXcoder-7B-v2 model’s technical report [10] and open-source repository [^28] https://github.com/aixcoder-plugin/aixcoder-7b-v2 only provide a prompt format for file-level completion and do not specify how to pass cross-file context.
-Therefore, we only evaluated this model in the file-level code completion tasks;
+Therefore, we only evaluated this model in the file-level code completion tasks.
 To conserve space in the data tables, we abbreviate Instruct to Inst and DeepSeek to DS.
@@
 Our implementation is based on the reproduction project of Qwen Coder [^30] https://github.com/QwenLM/Qwen3-Coder/tree/main/qwencoder-eval.
-We filtered out test cases where the total length of the context and the expected completion exceeded 8192 tokens (as tokenized by Qwen2.5-Coder-7B-Instruct) or where the generation length was greater than 64 tokens.
+We filtered out test cases where the total length of the context and the expected completion exceeded 8192 tokens or where the generation length was greater than 64 tokens.
 During evaluation, the temperature was set to 0, and the maximum output length of the model was capped at 64 tokens.
@@
 We use Pass@1 and Edit Similarity (ES) to measure model performance.
-For each model, we employed its corresponding prompt format for completion, with standardized limits of 4096 tokens for the intra-file context, 4096 tokens for the cross-file context, and a maximum output length of 160 tokens.
+For each model, we employed its corresponding prompt format for completion, with limits of 4096 tokens for the intra-file context, 4096 tokens for the cross-file context, and a maximum output length of 160 tokens.
 The temperature was set to 0.
@@
 TABLE VI: Suffix repetition ratio by model and dataset.
-Model SC* AX* ERE* CoLT
-StarCoder2-7B 0.33% 1.38% 30.58% 1.87%
-DS-Coder-6.7B-Inst 0.13% 1.62% 4.90% 1.98%
-DS-Coder-V2-Lite-Inst 0.15% 0.64% 2.15% 2.19%
-DS-Coder-6.7B-Base 0.13% 0.77% 4.73% 2.10%
-DS-Coder-V2-Lite-Base 0.15% 1.00% 4.64% 2.11%
-Qwen2.5-Coder-7B-Inst 0.13% 3.23% 1.03% 2.58%
-Qwen3-Coder-Flash 0.10% 1.16% 0.26% 1.62%
-Qwen2.5-Coder-7B 0.23% 0.69% 0.09% 1.51%
-Seed-Coder-8B-Inst 0.31% $\uparrow$210% 2.74% $\uparrow$621% 0.43% $\uparrow$377% 2.54% $\uparrow$418%
-Seed-Coder-8B-Base 0.17% $\uparrow$70% 1.67% $\uparrow$339% 2.23% $\uparrow$2377% 1.77% $\uparrow$261%
-SynthCoder-SFT 0.10% -0% 0.60% $\uparrow$57% 2.58% $\uparrow$2766% 1.42% $\uparrow$189%
-SynthCoder-DPO-wo-RS 0.11% $\uparrow$10% 0.52% $\uparrow$36% 0.60% $\uparrow$566% 0.50% $\uparrow$2%
-SynthCoder-DPO-RS 0.10% 0.38% 0.09% 0.49%
-    SC: SantaCoder; AX: aiXcoder; ERE: ExecRepoEval.
+Model AX* ERE* CoLT
+StarCoder2-7B 1.38% 30.58% 1.87%
+DS-Coder-6.7B-Inst 1.62% 4.90% 1.98%
+DS-Coder-V2-Lite-Inst 0.64% 2.15% 2.19%
+DS-Coder-6.7B-Base 0.77% 4.73% 2.10%
+DS-Coder-V2-Lite-Base 1.00% 4.64% 2.11%
+Qwen2.5-Coder-7B-Inst 3.23% 1.03% 2.58%
+Qwen3-Coder-Flash 1.16% 0.26% 1.62%
+Qwen2.5-Coder-7B 0.69% 0.09% 1.51%
+Seed-Coder-8B-Inst 2.74% $\uparrow$621% 0.43% $\uparrow$377% 2.54% $\uparrow$418%
+Seed-Coder-8B-Base 1.67% $\uparrow$339% 2.23% $\uparrow$2377% 1.77% $\uparrow$261%
+SynthCoder-SFT 0.60% $\uparrow$57% 2.58% $\uparrow$2766% 1.42% $\uparrow$189%
+SynthCoder-DPO-wo-RS 0.52% $\uparrow$36% 0.60% $\uparrow$566% 0.50% $\uparrow$2%
+SynthCoder-DPO-RS 0.38% 0.09% 0.49%
+    AX: aiXcoder; ERE: ExecRepoEval.
 TABLE VII: Prefix repetition ratio by model and dataset.
-Model SC* AX* ERE* CoLT
-StarCoder2-7B 0.00% 0.49% 0.17% 0.04%
-DS-Coder-6.7B-Inst 0.00% 0.43% 0.17% 0.03%
-DS-Coder-V2-Lite-Inst 0.00% 0.41% 0.60% 0.08%
-DS-Coder-6.7B-Base 0.00% 0.57% 0.34% 0.06%
-DS-Coder-V2-Lite-Base 0.00% 0.52% 0.95% 0.08%
-Qwen2.5-Coder-7B-Inst 0.00% 0.17% 0.43% 0.04%
-Qwen3-Coder-Flash 0.00% 0.15% 0.09% 0.04%
-Qwen2.5-Coder-7B 0.00% 0.18% 0.43% 0.04%
-Seed-Coder-8B-Inst 0.00% -0% 0.33% $\uparrow$560% 0.09% -0% 0.08% $\uparrow$300%
-Seed-Coder-8B-Base 0.00% -0% 0.56% $\uparrow$1020% 0.09% -0% 0.08% $\uparrow$300%
-SynthCoder-SFT 0.00% -0% 0.15% $\uparrow$200% 0.34% $\uparrow$278% 0.16% $\uparrow$700%
-SynthCoder-DPO-wo-RS 0.00% -0% 0.10% $\uparrow$100% 0.17% $\uparrow$89% 0.08% $\uparrow$300%
-SynthCoder-DPO-RS 0.00% 0.05% 0.09% 0.02%
-    SC: SantaCoder; AX: aiXcoder; ERE: ExecRepoEval.
+Model AX* ERE* CoLT
+StarCoder2-7B 0.49% 0.17% 0.04%
+DS-Coder-6.7B-Inst 0.43% 0.17% 0.03%
+DS-Coder-V2-Lite-Inst 0.41% 0.60% 0.08%
+DS-Coder-6.7B-Base 0.57% 0.34% 0.06%
+DS-Coder-V2-Lite-Base 0.52% 0.95% 0.08%
+Qwen2.5-Coder-7B-Inst 0.17% 0.43% 0.04%
+Qwen3-Coder-Flash 0.15% 0.09% 0.04%
+Qwen2.5-Coder-7B 0.18% 0.43% 0.04%
+Seed-Coder-8B-Inst 0.33% $\uparrow$560% 0.09% -0% 0.08% $\uparrow$300%
+Seed-Coder-8B-Base 0.56% $\uparrow$1020% 0.09% -0% 0.08% $\uparrow$300%
+SynthCoder-SFT 0.15% $\uparrow$200% 0.34% $\uparrow$278% 0.16% $\uparrow$700%
+SynthCoder-DPO-wo-RS 0.10% $\uparrow$100% 0.17% $\uparrow$89% 0.08% $\uparrow$300%
+SynthCoder-DPO-RS 0.05% 0.09% 0.02%
+    AX: aiXcoder; ERE: ExecRepoEval.
 We analyzed the frequency of erroneous suffix and prefix repetitions in the generated outputs across the aforementioned datasets, with the results presented in Table VI and Table VII, respectively.
@@
 While the prefix repetition rate of DeepSeek-Coder is generally lower than that of the Qwen-Coder series, the opposite is true for their suffix repetition rates.
-The Seed-Coder series of models demonstrates lower repetition rates compared to the other model families.
+Noteworthly, The Seed-Coder series of models demonstrates lower repetition rates compared to the other model families.
 We also noted that comparing the model before and after fine-tuning (i.e., Seed-Coder-8B-Base vs.
 SynthCoder-SFT), the repetition rates exhibit fluctuations of varying magnitudes and directions across different test sets.
-Interestingly, even without the inclusion of specialized repetition-suppression data, the context repetition rate of SynthCoder-DPO-wo-RS is lower than that of SynthCoder-SFT.
-This is because outputs containing repetitions that are generated and rejected during the sampling process are subsequently used as the dispreferred examples in the preference pairs for DPO training.
+Interestingly, even without the inclusion of specialized repetition-suppression data, the repetition rate of SynthCoder-DPO-wo-RS is lower than that of SynthCoder-SFT.
+This is because outputs containing repetitions that are generated and rejected during the sampling process are subsequently used as the negative examples in the preference pairs for DPO training.
 After incorporating an additional 11% of repetition-suppression data, SynthCoder-DPO-RS demonstrated a significantly greater reduction in this behavior.
@@
 These crucial capabilities are not adequately captured by current evaluation benchmarks.
-In summary, enhancing the performance of code completion necessitates a collaborative effort between LLM and client-side.
+In summary, enhancing the performance of code completion necessitates a collaborative effort between LLM and client.
 V-B About the Context Ranges

### v3

@@
 First, we fine-tune the model using Curriculum Learning technology.
-Following this, we perform alignment using Direct Preference Optimization (DPO) with preference pairs generated through Rejection Sampling.
+Following this, we perform alignment using Direct Preference Optimization (DPO) with preference pairs generated through rejected code sampling.
 Experimental results demonstrate that our final model excels on mainstream repository-level code completion benchmarks, including aiXcoder, ExecRepoBench, CrossCodeEval, and CoLT.
@@
 Examples include the GitHub Copilot [^1] https://github.com/features/copilot plugin, which is compatible with various IDEs, and the Cursor [^2] https://www.cursor.com IDE, which deeply integrates LLMs into various features of the development tool.
-These products are often built upon proprietary, cloud-based commercial models from providers like OpenAI and Anthropic, and have achieved remarkable success in both performance and reputation.
+These products are often built upon proprietary, cloud-based models from providers like OpenAI and Anthropic, and have achieved remarkable success in both performance and reputation.
 One study demonstrated that using GitHub Copilot can increase programming speed by 55.8% compared to a control group [4].
@@
   1. 1.
-    We propose a comprehensive methodology for code completion models that organically integrates and improves upon the data construction techniques, and training methods from several existing works [12, 13, 10], resulting in a new state-of-the-art (SOTA) model for FIM code completion.
+    We propose a comprehensive methodology for code completion models that organically integrates and improves upon the data construction and training methods from several existing works [12, 13, 10, 14], resulting in a new state-of-the-art (SOTA) model SynthCoder for FIM code completion based on Seed-Coder-8B-Base.
   2. 2.
-    We empirically evaluate several mainstream code completion models and discuss the gap between model capabilities and their successful application in completion tools, specifically addressing the issue of models erroneously repeating one or more lines of existing context.
+    We empirically evaluate several mainstream code completion models and discuss the gap between model capabilities and their successful application in completion tools, specifically addressing the issue of models erroneously repeating nearby context.
 The structure of this paper is as follows.
 Section III introduces the background and motivation of our research, and Section II reviews related work on applying LLMs to domains such as code completion.
-Section III provides a detailed account of our entire process, from data acquisition (Section III-A) and data processing (Section III-B) to training details (Section III-C).
+Section III provides a detailed account of our entire process, from data retrieval (Section III-A) and data construction (Section III-B) to details in model post-training(Section III-C).
 To validate the effectiveness of our method, in Section IV, we conduct a comprehensive evaluation of all baseline models and SynthCoder on file-level and repository-level code completion tasks, followed by an in-depth analysis of the results.
@@
 Section VI addresses the limitations of the current work.
-Finally, Section VII concludes the paper and suggests directions for future work.
+Finally, Section VII concludes the paper.
 II Related Work
-This section describes relevant work from the perspectives of code completion and adopting LLMs for code completion tasks.
+This section describes relevant work to position our work.
 II-A Fill in the Middle Code Completion
 Large Language Models (LLMs) for code have demonstrated significant potential in code completion tasks.
-Early works such as StarCoder and CodeLlama laid a foundational basis for this area [7, 6].
-However, constructing high-quality training data remains a challenge, particularly in ensuring that the completion target is meaningful in relation to its surrounding context (the prefix and suffix) and covers complex, real-world scenarios.
-Research in this domain has focused on structuring completion data to enhance model performance in authentic development contexts.
-Similarly leveraging code’s structural information, the ExecRepoBench team introduced a multi-level code completion method that also relies on ASTs to build training data.
-Their approach emphasizes enhancing the model’s capabilities in practical coding scenarios that involve complex, multi-file interdependencies [12].
-This reflects a trend in code completion research, which is progressively moving from the single-file level to more complex repository-level scenarios that include cross-file dependencies.
-II-B Repository Level Code Completion
+The Fill-in-the-Middle (FIM) paradigm, proposed by OpenAI [5], established a specialized input-output format for code completion that has since been widely adopted.
+Early works such as CodeLlama and StarCoder laid a foundational basis for this area [7, 6].
+However, constructing high-quality training data remains a challenge, particularly in ensuring that the completion target is meaningful in relation to its surrounding context and covers complex, real-world scenarios.
+Research in this domain has focused on structuring completion data to enhance model performance in authentic development contexts [11].
+Recent work has further explored syntax-aware FIM pretraining and evaluation.
+For instance, [15] proposed AST-T5, which incorporates AST structure into pretraining without architectural changes, showing strong performance in code-to-code tasks.
+[16] introduced SAFIM, a syntax-aware FIM benchmark that evaluates LLMs on structured code completions, revealing that pretraining methods and data quality often outweigh model size.
+[17] extended this line by proposing structure-aware FIM pretraining with syntax-tree-based masking, though their method focuses on single or sibling AST nodes and lacks open-source implementation.
+Other studies have addressed tokenization issues in FIM.
+[18] proposed a character-level infilling method to mitigate sub-token misalignment, which is relevant though not directly focused on syntax trees.
 Understanding the complex interdependencies among multiple files is crucial for handling modern software projects, a task known as Repository-Level Code Completion.
@@
 In terms of architectural innovation, the Code Graph Model (CGM) represents a repository as a heterogeneous graph (containing seven types of nodes and five types of edges) and integrates this structural information into the LLM’s attention layers via a graph-aware attention mask [24].
-II-C Post-training LLMs for Coding
+II-B Post-training LLMs for Coding
 Code generation is a significant application of LLMs.
-The core technology is founded on the Transformer architecture, which replaced Recurrent Neural Networks (RNNs) [20] and Convolutional Neural Networks (CNNs) [21].
-By leveraging a self-attention mechanism, Transformers effectively model long-range dependencies and exhibit excellent scaling stability, laying the groundwork for learning from code corpora [22].
-The decoder-only architecture has become the mainstream choice for code LLMs due to its advantages in text generation, including code.
 GPT-3 (175B) [25] demonstrated general-purpose capabilities (including code generation) without fine-tuning, and its emergent abilities [26] greatly spurred exploration in the field.
 Subsequently, OpenAI trained the CodeX model on a vast amount of open-source code (e.g., based on GPT-3.5), revealing for the first time the broad prospects of LLMs in software engineering.
-The open-source community has since produced powerful base models, such as Llama/Llama2 [25, 26], the Qwen series, and the DeepSeek series, as well as code-specific models like DeepSeek-Coder [27], Qwen2.5-Coder [8], and Seed-Coder [9].
-The base model used for fine-tuning in this paper is from the Seed-Coder series.
+The open-source community has since produced powerful base models specifically for code-related tasks, like the DeepSeek-Coder [27], Qwen2.5-Coder [8], and Seed-Coder [9] series.
 These models have achieved outstanding results on code benchmarks, with some metrics even surpassing those of larger, closed-source models.
@@
 Although reasoning models are constrained in low-latency tasks like code completion, they are becoming indispensable for more systematic software engineering tasks such as code repair, repository-level code understanding, and complex code generation.
+While the field of code generation is evolving towards autonomous agents—benchmarked by SWE-bench [30] and embodied by frameworks like SWE-agent [31] and OpenDevin [32]—their iterative, tool-heavy nature induces significant latency.
+This makes them unsuitable for real-time code completion.
 Another critical path to enhancing code LLM performance is the construction of high-quality, domain-specific training data.
@@
 To address this, subsequent research proposed innovative methods: Evol-Instruct [34] evolutionarily rewrites simple instructions, while OSS-Instruct [35] innovatively uses open-source code as seeds to prompt an LLM to generate relevant problems and solutions.
-In summary, current research on code Large Language Models focuses on three key directions: (1) Constructing high-quality Fill in the Middle (FIM) training data by leveraging structural information like Abstract Syntax Trees (ASTs) to enhance the model’s code completion capabilities in both single-file and complex cross-file dependency scenarios [11, 12]; (2) Advancing the code completion task from the single-file level to the Repository Level, which requires an understanding of the overall project structure and cross-file dependencies.
-Relevant research encompasses innovative paradigms such as retrieval augmentation [15, 16], graph structure modeling [17, 15], and the integration of static analysis [33, 19]; (3) Continuously improving model performance on complex tasks like code generation, understanding, and repair by combining advanced post-training techniques – especially diverse instruction data synthesis methods [31, 32] – with innovative model architectures and optimization strategies [28].
-Collectively, these research streams lay the foundation for building more powerful and practical intelligent code assistants.
-Unfortunately, due to inherent constraints in their internal mechanisms, these different approaches cannot be simply combined to boost the performance of code completion models [34].
+Meanwhile, [36] aligned FIM training with code generation tasks, demonstrating the versatility of FIM-based training paradigms beyond completion.
 III Methodology
-The details in training SynthCoder will be elaborated in this section.
+The details in the methodology SynthCoder will be elaborated in this section.
 III-A Data Retrieval
-To build a high-quality and multi-dimensional training corpus for code completion, we constructed data for different application scenarios on two levels.
-First, we utilized The-Stack-V2 [35] open-source repository dataset to create Fill-in-the-Middle (FIM) code completion samples that include both intra-file and cross-file contexts by masking abstract syntax tree nodes.
-Second, based on the pre-training corpus from the OpenCoder project, we generated samples for completing function bodies based on function signatures and docstrings.
-Finally, to improve training efficiency and model performance, we applied a unified quality filtering process to all data, which combined a heuristic rule pipeline with sample selection based on model perplexity.
-III-A1 Files from Open-Source Repositories
-To construct a code completion dataset with both intra-file and cross-file contexts, with considerations of data quality and the costs of data acquisition, we extracted files in Java, Python, C++, Go, JavaScript, and TypeScript from repositories with 256 or more stars in the the-stack-v2-train-smol-ids [^6] https://huggingface.co/datasets/bigcode/the-stack-v2-train-smol-ids, a subset of The-Stack-V2, which served as the basis for our AST node masking and cross-file context construction.
-III-A2 Snippets of Standalone Functions
-For the scenario of completing function bodies based on function signatures and docstrings, we used the pre-training annealing data, opc-annealing-corpus [^7] https://huggingface.co/datasets/OpenCoder-LLM/opc-annealing-corpus/, from the OpenCoder project to construct function completion samples.
-This dataset contains over ten million lines of code snippets synthesized by GPT-4.
-Specifically, we extracted function implementations from complete code files to serve as the code snippets to be completed, while the function signature, docstring, and other external code related to the target function collectively formed the context.
-For Java, Python, C++, Go, JavaScript, and TypeScript, we used data from the synthetic_code_snippet subset.
-As the Go language was not included in this subset, we used the synthetic_qa subset for extraction instead.
-Since functions in this dataset may lack docstrings, we used the DeepSeek-V3-0324888https://huggingface.co/deepseek-ai/DeepSeek-V3-0324 model to synthesize corresponding docstrings for the extracted functions written in Go.
-III-A3 Filtering Based on Data Quality
-To improve training efficiency and enhance the final model’s performance, all data from our sources underwent a quality filtering process.
-OpenCoder Pipeline
-OpenCoder Data Filtering Pipeline is a heuristic filtering framework for large-scale code pre-training corpora, which considers the unique characteristics of different programming languages and includes over 100 filtering rules.
-The guiding principle of this pipeline is to remove harmful data while ensuring that the overall distribution of the dataset is not significantly affected [13].
-In this paper, we adopt this pipeline [^9] https://github.com/OpenCoder-llm/opc_data_filtering to filter the raw code data based on heuristic rules, thereby obtaining high-quality yet diverse code, which in turn improves the overall quality of the subsequently synthesized code completion data.
-Perplexity Filtering
+We started with the an officially deduplicated subset [^6] https://huggingface.co/datasets/bigcode/the-stack-v2-train-smol-ids of Stack v2 [37] dataset, which is one of the biggest code corpora.
+To balance data quality and acquisition cost, we curated a specialized subset by selecting files from repositories with 256 or more stars, focusing on six languages: Java (25%), Python (25%), Go (14%), JavaScript (14%), C++ (14%), and TypeScript (8%).
+This final collection served as the basis for our Fill-in-the-Middle (FIM) training set construction.
+The raw data was subsequently filtered through a two-stage filtering process: a heuristic rule-based pipeline and a model perplexity-based sample selection.
+III-A1 Filtering Based on Heuristics
+In this work, we also OpenCoder’s opc_data_filtering [^7] https://github.com/OpenCoder-llm/opc_data_filtering, which is a heuristic filtering framework for large-scale code pre-training corpora, which considers the unique characteristics of different programming languages and includes over 100 filtering rules [13], with guiding principle to remove harmful data while ensuring that the overall distribution of the dataset is not significantly affected, thus obtaining a code dataset with relatively high quality yet remain diversity.
+III-A2 Filtering Based on Perplexity
 Perplexity (PPL) measures the certainty of a model’s predictions for a given sample.
@@
 III-B Data Construction
-The core of our data construction lies in deeply simulating the real-world coding habits and contextual needs of developers.
-Our method encompasses three key aspects: (1) For constructing the completion target, we combine fine-grained node extraction based on Abstract Syntax Trees (ASTs) with heuristic strategies that mimic user behavior.
-(2) For building the completion context, we introduce critical cross-file information using code similarity (BM25) and code dependency graphs (Call Graphs).
-(3) For optimizing data quality, we employ rejection sampling and a multi-stage filtering process to generate high-quality preference pairs for alignment training, aiming to enhance the model’s FIM capabilities and suppress context repetition.
-Through this multi-dimensional construction strategy, we aim to generate high-quality training data that more closely resembles real-world application scenarios.
+The core objective of our dataset construction is to simulate, to the greatest extent possible, the coding practices, contextual needs, and code completion expectations of human engineers in real-world file-level and repository-level programming tasks.
+This section elaborates the process to construct such a training corpora to meet this objective.
 III-B1 Inner File Context
@@
 After breaking down complex features, developers can concentrate on implementing smaller, logically simpler code segments.
-This method of progressively decomposing a complex system naturally aligns with how parsing tools break code down into a tree-like structure.
 In this paper, we use the Tree-sitter [^8] https://tree-sitter.github.io/tree-sitter parsing tool to parse code files into abstract syntax trees.
@@
     • Syntax-Aware Triggered Completion: The completion is triggered immediately following a language-specific, syntactically significant token.
-We define a curated set of such tokens for each programming language, encompassing critical keywords (e.g., for control flow, definitions) and operators (e.g., for assignment, member access).
-The remainder of the line after the token is designated as the completion target.
-  • Code Fragments Within Parentheses (4.86%): Paired parentheses often enclose relatively independent code snippets.
-In code editors, typing an opening parenthesis automatically generates the closing one and places the cursor in between.
+We define a curated set of such tokens for each programming language, encompassing critical keywords (e.g., for control flow, definitions) and operators (e.g., for assignment, member access), which the remainder of the line after these token will be designated as the completion target.
+  • Code Fragments Within Parentheses (4.86%): In code editors, typing an opening parenthesis automatically generates the closing one and places the cursor in between.
 We sample data to replicate this common scenario.
-  • Code Fragments After Comments (2.92%): When a model cannot directly understand a developer’s intent, the developer may manually write a comment describing the functionality of the code block to be written, then trigger code completion on the next line.
-To simulate this, we use Tree-sitter to detect full-line comment nodes in code files and treat the first code block immediately following the comment as the completion target.
+  • Code Fragments After Comments (2.92%): The developer may manually write a comment describing the code block to be written, then trigger completion on the next line.
+To simulate this, we detect full-line comment nodes in code files and treat the first code block immediately following the comment as the completion target.
 Strategies Based on Random Single or Multiple Lines
@@
 The query consists of the code to be completed along with some of its adjacent context.
-We then compute the BM25 similarity score between this query and all code chunks from other files within the project, selecting the highest-scoring snippets (sum up within a maximum total length after tokenization) to serve as cross-file context.
+We then compute the BM25 similarity score between this query and all code chunks from other files within the project, selecting the highest-scoring snippets (sum up within a maximum total length after tokenization) to serve as the retrieval based-part of cross-file context.
 Dependency Related Code Snippets
 Code import statements directly indicate which external APIs and modules can be called within a file.
-Retrieving and providing such code information from the project allows the LLM to better decide when to use imported modules and how to use them more correctly.
+Retrieving and providing such code information from the project allows the LLM to better decide when to and how to use the imported snippets correctly.
 This paper uses the open-source Repo-specific Semantic Graph (RSSG) call graph construction module from the RepoFuse project [43] to directly build call relationships between different files from the project structures of open-source repositories collected from The-Stack-V2 dataset.
@@
 These scripts extract class definitions, method signatures, functional description comments, and other key information from the called code files, thereby significantly reducing the context length while retaining essential information.
-III-B3 Rejection Sampling
+III-B3 Rejected Code Sampling
 To further optimize the model’s FIM capabilities, this paper introduces paired preference alignment training.
@@
 The main steps are as follows:
-Rejection Sampling Strategies
-Based on a cross-file code dataset constructed using the BM25 algorithm and call relationship graphs, we use the target model intended for alignment on each sample.
-With the sampling temperature set to 1.0, we generate 10 candidate negative samples via rejection sampling [10].
+Rejected Code Sampling Strategies
+Based on the dataset containing cross-file information, we use the best checkpoint in the finetuning stage which is the model intended for alignment to generate 10 candidate negative samples via rejected code sampling [10] on each sample with temperature set to 1.0.
 Multi-stage Filtering Strategies
@@
 Subsequently, similarity filtering based on BLEU scores is used to remove negative samples that are too similar to the ground truth.
+After the screening process, we retain up to 3 negative samples for each original data point, which constitute the paired dataset for preference alignment training.
 Contextual Repetition Suppression
-After the screening process, we retain up to 3 negative samples for each original data point, which constitute the paired dataset for preference alignment training.
-At the same time, to suppress the model’s tendency to simply repeat the context, we randomly sampled 10% of the data from rejection sampling where the expected completion does not start with the first line of the suffix.
-For this subset, we constructed suffix-repetition suppression pairs where the positive example is the ground-truth completion and the negative example is the first line of the suffix code.
+To suppress the model’s tendency to simply repeat the context, we randomly sampled 10% of the data from rejected code sampling where the expected completion does not start with the first line of the suffix.
+We constructed suffix-repetition suppression pairs where the positive example is the ground-truth completion and the negative example is the first line of the suffix code.
 This data was then added to the synthesized dataset.
@@
 III-C Model Post-training
-Using the data synthesized with the methods described above, we conducted two stages of post-training on the Seed-Coder-8B-Base model [9]: fine-tuning and alignment.
-Seed-Coder-8B-Base is an 8B-parameter open-source large language model for code developed by the ByteDance Seed team.
-It serves as the base model for the Seed-Coder series and is designed to provide a powerful starting point for various downstream tasks, including code generation and reasoning.
-Thus, we use Seed-Coder-8B-Base as our based model for code completetion task.
+During both training stages, the SPM (Suffix-Prefix-Middle) FIM prompt format is adopted, as described in its technical report [9]. In reference of [17], we set FIM rate to 0.7.
 III-C1 Finetune Stage
@@
 Curriculum Learning
-The data preprocessing for curriculum learning fine-tuning is primarily divided into three steps:
-  • Step 1: Code Parsing.
-    After initial filtering, we use Tree-sitter to parse code files and generate Abstract Syntax Trees (ASTs).
-  • Step 2: Complexity Sorting.
-    We measure complexity by the number of identifier nodes in the AST and sort them in descending order. Symbols are extracted by recursively traversing the AST.
-  • Step 3: Sample Selection.
+The data preprocessing for curriculum learning fine-tuning is divided into two steps:
+  • Step 1: Complexity Sorting.
+    We use Tree-sitter to parse code files and generate Abstract Syntax Trees (ASTs), then measure complexity by the number of identifier nodes in the AST and sort them in descending order as is described in [14].
+Symbols are extracted by recursively traversing the ASTs.
+  • Step 2: Sample Selection.
     We select the top $k\%$ (k=30) most complex fragments to serve as FIM and completion training samples, respectively, forming the final training dataset.
 Training Details
-In the curriculum learning fine-tuning stage, we trained the model for one epoch using the AdamW optimizer, with a learning rate set to 3e-6 combined with a 50-step warmup.
-The training data selected was the top 30% (k=0.3) of the original dataset sorted by complexity, and the input sequences were uniformly truncated to within 8000 tokens.
-The entire process occupied 4 H800 80G PCIE GPUs, with a global batch size of 64 (micro-batch size of 1).
-Strategies such as DeepSpeed ZeRO-1, Flash Attention, and gradient checkpointing activation re-computation were used to improve training efficiency.
-Finally, the process accounted for 800 steps and finished in 20 hours.
+In the curriculum learning fine-tuning stage, we performed full-parameter fine-tuning on the Seed-Coder-8B-Base model, updating all of its weights.
+The model was trained for one epoch using the AdamW optimizer with a learning rate of $3\times 10^{-7}$ and a 50-step linear warmup.
+The training data consisted of the top 30% ($k=0.3$) most complex samples from the original dataset, with all input sequences truncated to a maximum length of 8192 tokens.
+The training was conducted on four NVIDIA H800 (80GB PCIe) GPUs, using a global batch size of 256 and a micro-batch size of 1 per GPU.
+To enhance efficiency, we employed several optimization strategies, including DeepSpeed ZeRO-1 [44], Flash Attention [45], and gradient checkpointing [^9] https://github.com/cybertronai/gradient-checkpointing.
+This fine-tuning process was completed in 800 steps over approximately 20 hours.
+Checkpoints are recorded every 100 steps and only the best checkpoint will be retained.
 III-C2 Alignment Stage
@@
 In the preference alignment stage using the DPO algorithm, the DPO parameter $\beta$ was set to 0.9, with an equivalent batch size of 64.
-The model was trained for one epoch with a maximum learning rate of 5e-5, a linear learning rate warmup of 48 steps, and a cosine learning rate decay scheduler.
-The maximum length for training data was 8192.
-For LoRA, the rank and alpha were set to 32 and 64, respectively, and the LoRA training modules included various modules within Self-Attention and the feed-forward networks.
-The training was conducted on 4 NVIDIA H800 80G PCIE GPUs using DeepSpeed ZeRO-2 and a CPU offloading strategy.
+The model was trained for one epoch with a maximum learning rate of 1e-4, a linear learning rate warmup of 32 steps, and a cosine learning rate decay scheduler.
+The maximum length for each training data was 8192 tokens.
+For parameter-efficient fine-tuning, we employed the Low-Rank Adaptation (LoRA) technique [47].
+The LoRA rankrank was set to 32 and the scaling factor alphaalpha was set to 64.
+The training was conducted on 4 NVIDIA H800 80G PCIE GPUs, leveraging the TRL framework [^10] https://github.com/huggingface/trl (version 0.17.0) in conjunction with DeepSpeed ZeRO-2 [44] (with a CPU offloading strategy) and gradient checkpointing.
 The changes in Loss, Reward Accuracy, and Reward Margin during the training process are shown in Figure 2.
 The alignment process lasted for 400 steps within 24 hours, and it can be observed that the Training Loss decreased relatively steadily, while the Reward Accuracy and Reward Margin increased and then remained relatively stable.
+Checkpoints are recorded every 100 steps and only the best checkpoint will be retained.
 (a) Training loss of DPO.
@@
 We conducted a comprehensive and rigorous evaluation of SynthCoder and baseline models on both file-level and repository-level code completion tasks.
-Due to constraints in computational resources, we selected representative open-source models with Fill-in-the-Middle (FIM) capabilities that were released between 2024 and the first half of 2025.
+IV-A Evaluation Setup
+IV-A1 Benchmark Baselines
+Due to constraints in computational resources, we selected representative open-source models with Fill-in-the-Middle (FIM) capabilities with size ranging from 6B to 8B that were released between 2024 and the first half of 2025.
 Models released earlier, such as CodeLlama-7b-hf [^11] https://huggingface.co/codellama/CodeLlama-7b-hf from 2023, were excluded.
-Additionally, we excluded models not specifically fine-tuned for FIM tasks, such as MagicCoder [^12] https://huggingface.co/ise-uiuc/Magicoder-S-DS-6.7B.
-All models were deployed on a single local H800 80G PCIe GPU, and inference was performed using VLLM version 0.9.1131313https://github.com/vllm-project/vllm.
 The final set of models used for comparison is as follows.
-  • aiXcoder-7b-v2141414https://huggingface.co/aiXcoder/aixcoder-7b-v2
-    A 7B parameter code LLM designed to fully leverage long-context information, particularly excelling at cross-file code completion.
-  • StarCoder2-7B [^15] https://huggingface.co/bigcode/starcoder2-7b
-    A 7B parameter code LLM developed by the BigCode project. It was trained on The-Stack-V2, a large-scale dataset containing multiple programming languages.
-  • DeepSeek-Coder-6.7B (-base [^16] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base, -instruct [^17] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct)
-    A pair of 6.7B code LLMs developed by DeepSeek. The former is the base model, while deepseek-coder-6.7b-instruct is the instruction-tuned version of deepseek-coder-6.7b-base.
-  • Qwen2.5-Coder-7B [^18] https://huggingface.co/Qwen/Qwen2.5-Coder-7B (-Instruct [^19] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct)
-    A pair of 7B code LLMs developed by Alibaba, designed for code understanding and generation.
-Qwen2.5-Coder-7B is the base model, while Qwen2.5-Coder-7B-Instruct is the instruction-tuned version of Qwen2.5-Coder-7B.
-  • Seed-Coder-8B (-Base [^20] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Base, -Instruct [^21] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Instruct)
-    A pair of 8B parameter code LLMs developed by the ByteDance.
-The model was trained using a model-driven data selection method, aiming to efficiently enhance its code generation abilities, while Seed-Coder-8B-Instruct is the instruction-tuned version of Seed-Coder-8B-Base.
+  • aiXcoder-7B-V2121212https://huggingface.co/aiXcoder/aiXcoder-7B-V2 A 7B parameter code LLM designed to fully leverage long-context information, particularly excelling at cross-file code completion.
+  • StarCoder2-7B [^13] https://huggingface.co/bigcode/starcoder2-7b A 7B parameter code LLM developed by the BigCode project.
+  • DeepSeek-Coder-6.7B (-base [^14] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base, -instruct [^15] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct) A pair of 6.7B code LLMs developed by DeepSeek, while deepseek-coder-6.7b-instruct is the instruction-tuned version of deepseek-coder-6.7b-base.
+  • Qwen2.5-Coder-7B [^16] https://huggingface.co/Qwen/Qwen2.5-Coder-7B (-Instruct [^17] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct) A pair of 7B code LLMs developed by Alibaba, while Qwen2.5-Coder-7B-Instruct is the instruction-tuned version.
+  • Seed-Coder-8B (-Base [^18] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Base, -Instruct [^19] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Instruct) A pair of 8B parameter code LLMs developed by the ByteDance, while Seed-Coder-8B-Instruct is the instruction-tuned version of Seed-Coder-8B-Base.
+Additionally, we include Mixture-of-Experts (MoE) models that, despite having fewer than 10B active parameters, possess a total parameter count that exceeds this threshold.
   • DeepSeek-Coder-V2-Lite (-Base [^20] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Base, -Instruct [^21] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct)
-    A pair of lightweight code LLMs in DeepSeek-Coder-V2 model family [^24] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Base.
-Both of them utilize a Mixture-of-Experts (MoE) architecture, while DeepSeek-Coder-V2-Lite-Instruct is the instruction-tuned version of DeepSeek-Coder-V2-Lite-Base.
+    A pair of lightweight code LLMs in DeepSeek-Coder-V2 model family [^22] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Base, while DeepSeek-Coder-V2-Lite-Instruct is the instruction-tuned version.
   • Qwen3-Coder-30B-A3B-Instruct [^23] https://huggingface.co/Qwen/Qwen3-Coder-30B-A3B-Instruct
-    An efficient instruction-tuned code model from Alibaba’s Qwen3 series, which also employs a Mixture-of-Experts (MoE) architecture and possesses excellent instruction-following capabilities.
+    An efficient instruction-tuned code model from Alibaba’s Qwen3 Coder series.
 To ensure the rigor of our experimental results, we only included models in our comparison that were evaluated in the aforementioned file-level FIM code completion benchmarks and for which the official developers provided a specific FIM prompt format in their technical reports or open-source repository documentation.
-  • The Qwen-Coder series (including Qwen2.5-Coder and Qwen3-Coder) models follow the format specified in their technical report [8];
-  • The DeepSeek-Coder series models follow the example provided in the official repository [^26] https://github.com/deepseek-ai/DeepSeek-Coder;
-  • The Seed-Coder series models follow the official response in an issue on their official open-source repository [^27] https://github.com/ByteDance-Seed/Seed-Coder/issues/4;
-  • The StarCoder2 model follows the format described in its technical report [35];
-  • The aiXcoder-7B-v2 model’s technical report [10] and open-source repository [^28] https://github.com/aixcoder-plugin/aixcoder-7b-v2 only provide a prompt format for file-level completion and do not specify how to pass cross-file context.
-Therefore, we only evaluated this model in the file-level code completion tasks.
-To conserve space in the data tables, we abbreviate Instruct to Inst and DeepSeek to DS.
-To validate the impact of different training stages and data compositions, we evaluated three versions of our model: SynthCoder-8B-SFT, SynthCoder-8B-DPO-wo-RS, and SynthCoder-8B-DPO-RS (the final version, referred to as SynthCoder-8B).
+Note that the aiXcoder-7B-V2 model’s technical report [10] and open-source repository [^24] https://github.com/aixcoder-plugin/aiXcoder-7B-V2 only provide a prompt format for file-level completion and do not specify how to pass cross-file context.
+Therefore, we only evaluated it in the file-level code completion tasks.
+IV-A2 SynthCoder Variants for Ablation Study
+To validate the impact of different training stages and data compositions, we evaluated three versions of our model: SynthCoder-8B-SFT, SynthCoder-8B-DPO-wo-RS, and SynthCoder-8B-DPO (the final version, referred to as SynthCoder-8B).
 We compared these against the base model, Seed-Coder-8B-Base.
 The differences between these model versions are detailed in Table I.
-To save space, all numerical values in tables are rounded to one decimal place.
-The percentage of increase or decrease refers to the change relative to SynthCoder-8B-DPO-RS.
 TABLE I: Comparisons between SynthCoder and its varients.
@@
 SynthCoder-8B-DPO-wo-RS YES YES NO
-SynthCoder-8B-DPO-RS YES YES YES
-IV-A File Level Completion
-We select aiXcoder dataset as the evaluation dataset to verify our model’s capacity on file level code completetion.
-We utilized the dataset [^29] https://github.com/aixcoder-plugin/aiXcoder-7B/tree/main/aiXcoder_ex-tended_dataset provided by [11] to evaluate the model’s Fill-in-the-Middle (FIM) code completion performance on files from real-world open-source projects.
-The specific evaluation metrics used were Exact Match (EM) and Edit Similarity (ES).
-Our implementation is based on the reproduction project of Qwen Coder [^30] https://github.com/QwenLM/Qwen3-Coder/tree/main/qwencoder-eval.
-We filtered out test cases where the total length of the context and the expected completion exceeded 8192 tokens or where the generation length was greater than 64 tokens.
-During evaluation, the temperature was set to 0, and the maximum output length of the model was capped at 64 tokens.
-As shown in Table II, SynthCoder-8B achieves an improvement of over 10% after finetuning, surpassing all baseline models, including those with a larger total parameter count such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B.
-Specifically, the fine-tuned SynthCoder-8B-SFT demonstrates an overall improvement of more than 10% compared to the base model, with a maximum gain of 22% on the JavaScript EM score.
-The aligned SynthCoder-8B-DPO-wo-RS shows further modest gains across all metrics, improving upon SynthCoder-8B-SFT by 4.2% in overall EM and 1.6% in overall ES.
-The final model with repetition suppression, SynthCoder-8B-DPO-RS, trades a slight decrease of up to 1.2% in C++ ES and on Python metrics for a 2.1% increase in overall EM when compared to SynthCoder-8B-DPO-wo-RS.
-Ultimately, this results in a total improvement of 17.8% in overall EM and 12.3% in overall ES relative to the base model.
+SynthCoder-8B-DPO YES YES YES
+IV-A3 Benchmark Datasets
+We select aiXcoder dataset as the evaluation dataset to verify our model’s capacity on file level code completetion, and ExecRepoBench, CrossCodeEval, and CoLT to evaluate the LLM’s capabilities in repository-level code completion, as is shown below.
+There’s always been only one trial for each model in all the benchmarks respectively.
+All models were deployed on a single local H800 80G PCIe GPU, and inference was performed using VLLM version 0.9.1252525https://github.com/vllm-project/vllm.
+For all the generation process, we employed deterministic decoding with the following hyperparameters: $temperature=0$, top_p=1.0top\_p=1.0, frequency_penalty=0.0frequency\_penalty=0.0, repetition_penalty=1.0repetition\_penalty=1.0, and length_penalty=1.0length\_penalty=1.0.
+aiXcoder
+The aiXcoder dataset [11] evaluates Fill-in-the-Middle (FIM) code completion on real-world open-source projects.
+Our evaluation follows the Qwen Coder reproduction script [^26] https://github.com/QwenLM/Qwen3-Coder/tree/main/qwencoder-eval.
+We filtered out test cases where the combined length of the context and ground truth exceeded 8192 tokens, or where the ground truth itself was longer than 64 tokens.
+During inference, we used a temperature of 0 and capped the maximum generation length at 64 tokens.
+CrossCodeEval with BM25 Retrieval
+CrossCodeEval [19] is a benchmark designed to assess cross-file code completion.
+It uses static analysis to create test cases from projects in Python, Java, TypeScript, and C# that are only solvable with external context.
+For our evaluation, we use the officially provided context retrieved via the BM25 algorithm.
+This context includes the top 5 most relevant code snippets, each truncated to a maximum of 10 lines.
+We adhere to each model’s specific prompt format, setting the maximum input length to 8192 tokens and the maximum output to 50 tokens.
+CoLT with Dependency-based Retrieval
+The CoLT benchmark [10] assesses cross-file context utilization across 12,000 samples in Python, Java, C/C++, and Go.
+To specifically evaluate the model’s ability to leverage code structure, we exclusively used the context retrieved through dependency call relationships.
+For each model, we followed its native prompt format, standardizing the maximum input length to 4096 tokens and the maximum output length to 200 tokens.
+ExecRepoBench
+ExecRepoBench [^27] https://huggingface.co/datasets/CSJianYang/ExecRepoBench evaluates code completion within real-world Python repositories [12].
+Uniquely, it provides test cases with both local, intra-file context in a Fill-in-the-Middle (FIM) format and relevant cross-file context as complete files.
+In our evaluation, we implemented a rule-based post-processing step to correct indentation errors in the generated code.
+The intra-file and cross-file contexts were each limited to 4096 tokens, and the maximum generation length was set to 160 tokens.
+IV-B Precision in Benchmarks
+The precision of our model on the benchmark datasets is reported in this section. The evaluation is based on the following metrics.
+  • Exact Match (EM): The model generation result is identical with the canonical solution. Adopted in aiXcoder, CrossCodeEval and CoLT.
+  • Edit Similarity (ES): The edit similarity between model generation result and canonical solution calculated by the FuzzyWuzzy library  [^28] https://github.com/seatgeek/fuzzywuzzy.
+Adopted in aiXcoder, CrossCodeEval, CoLT and ExecRepoBench.
+  • Edit Similarity RepoEval (ES_R): The edit similarity between model generation result and canonical solution calculated in the way of RepoEval [20]. Adopted only in CrossCodeEval.
+  • Pass@1: Whether the complete code assembled by the prefix, suffix and generation results can pass all the corresponding test cases. Adopted only in ExecRepoEval.
+To conserve space in the data tables, we abbreviate Instruct to Inst and DeepSeek to DS.
+To save space, when presenting the precision metrics, all numerical values in tables are rounded to one decimal place, and all the percentages are calculated between current varient of SynthCoder and the final SynthCoder-8B-DPO while blue means that the latter is better.
+IV-B1 aiXcoder
+As shown in Table II, SynthCoder-8B-SFT achieves an improvement of over 10% after finetuning, surpassing all baseline models, including those with a larger total parameter count such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B.
+Specifically, the fine-tuned SynthCoder-8B-SFT demonstrates an overall improvement of more than 10% compared to the base model, with a maximum gain of 20.3% on the JavaScript EM score.
+The aligned SynthCoder-8B-DPO-wo-RS shows further modest gains across all metrics, improving upon SynthCoder-8B-SFT by 4.1% in overall EM and 1.6% in overall ES.
+The final model with repetition suppression, SynthCoder-8B-DPO, trades a slight decrease of up to 1.2% in Python ES for a 2.1% increase in overall EM when compared to SynthCoder-8B-DPO-wo-RS.
+Ultimately, this results in a total improvement of 17.8% in overall EM and 12.3% in overall ES relative to the base model, showing the effectiveness of our approach.
 TABLE II: Performance of models on aiXcoder dataset [11].
@@
 Qwen2.5-Coder-7B 49.3 79.9 44.1 77.6 53.6 82.8 45.0 78.3 48.5 80.0
-Seed-Coder-8B-Inst 38.6 $\downarrow$28.2% 72.0 $\downarrow$12.8% 33.1 $\downarrow$37.4% 67.2 $\downarrow$17.2% 44.0 $\downarrow$28.4% 74.5 $\downarrow$15.3% 38.6 $\downarrow$16.9% 70.5 $\downarrow$11.2% 39.0 $\downarrow$27.6% 71.3 $\downarrow$14.3%
-Seed-Coder-8B-Base 44.9 $\downarrow$10.2% 75.3 $\downarrow$7.8% 35.0 $\downarrow$30.0% 66.9 $\downarrow$17.8% 46.7 $\downarrow$20.9% 75.3 $\downarrow$14.1% 41.6 $\downarrow$8.4% 72.5 $\downarrow$8.3% 42.3 $\downarrow$17.8% 72.6 $\downarrow$12.3%
-SynthCoder-8B-SFT 46.3 $\downarrow$6.7% 79.3 $\downarrow$2.4% 42.1 $\downarrow$8.0% 77.5 $\downarrow$1.6% 52.8 $\downarrow$6.9% 83.9 $\downarrow$2.4% 43.7 $\downarrow$3.2% 78.3 $\downarrow$0.1% 46.9 $\downarrow$6.3% 80.2 $\downarrow$1.7%
-SynthCoder-8B-DPO-wo-RS 48.8 $\downarrow$1.4% 81.4 $\uparrow$0.3% 43.5 $\downarrow$4.7% 78.5 $\downarrow$0.4% 55.1 $\downarrow$2.4% 85.2 $\downarrow$0.8% 45.2 $\uparrow$0.2% 79.4 $\uparrow$1.2% 48.8 $\downarrow$2.1% 81.5 $\downarrow$0.1%
-SynthCoder-8B-DPO-RS 49.5 81.2 45.5 78.8 56.5 85.9 45.1 78.5 49.8 81.5
-IV-B Repository Level Code Completion
-We selected ExecRepoBench, CrossCodeEval, and CoLT to evaluate the LLM’s capabilities in repository-level code completion.
-IV-B1 CrossCodeEval with BM25 Retrieval Strategy
-CrossCodeEval is a benchmark designed to evaluate the ability of LLMs to leverage cross-file context for code completion.
-Its data is sourced from open-source GitHub projects and covers four programming languages: Python, Java, TypeScript, and C#.
-The benchmark employs static analysis methods to construct test cases that are solvable only with the aid of cross-file context [14].
-In this paper, we adopt the same evaluation metrics as the original paper – EM, ES, and ES_repoeval (abbreviated as ES_R) – and use the official cross-file context data provided by CrossCodeEval, which is retrieved using the BM25 text similarity algorithm.
-Each retrieved code snippet is limited to a maximum of 10 lines, and the top 5 most relevant snippets are selected.
-For each model, we utilize its specific prompt format for completion, with a input context length of 8192 tokens, a maximum output length of 50 tokens, and a temperature of 0.
-As shown in Table III, SynthCoder-8B demonstrates a modest overall improvement compared to its base model and achieves the best performance among all models tested.
+Seed-Coder-8B-Inst 38.6 $\uparrow$28.2% 72.0 $\uparrow$12.8% 33.1 $\uparrow$37.4% 67.2 $\uparrow$17.2% 44.0 $\uparrow$28.4% 74.5 $\uparrow$15.3% 38.6 $\uparrow$16.9% 70.5 $\uparrow$11.2% 39.0 $\uparrow$27.6% 71.3 $\uparrow$14.3%
+Seed-Coder-8B-Base 44.9 $\uparrow$10.2% 75.3 $\uparrow$7.8% 35.0 $\uparrow$30.0% 66.9 $\uparrow$17.8% 46.7 $\uparrow$20.9% 75.3 $\uparrow$14.1% 41.6 $\uparrow$8.4% 72.5 $\uparrow$8.3% 42.3 $\uparrow$17.8% 72.6 $\uparrow$12.3%
+SynthCoder-8B-SFT 46.3 $\uparrow$6.7% 79.3 $\uparrow$2.4% 42.1 $\uparrow$8.0% 77.5 $\uparrow$1.6% 52.8 $\uparrow$6.9% 83.9 $\uparrow$2.4% 43.7 $\uparrow$3.2% 78.3 $\uparrow$0.1% 46.9 $\uparrow$6.3% 80.2 $\uparrow$1.7%
+SynthCoder-8B-DPO-wo-RS 48.8 $\uparrow$1.4% 81.4 $\downarrow$0.3% 43.5 $\uparrow$4.7% 78.5 $\uparrow$0.4% 55.1 $\uparrow$2.4% 85.2 $\uparrow$0.8% 45.2 $\downarrow$0.2% 79.4 $\downarrow$1.2% 48.8 $\uparrow$2.1% 81.5 $\uparrow$0.1%
+SynthCoder-8B-DPO 49.5 81.2 45.5 78.8 56.5 85.9 45.1 78.5 49.8 81.5
+IV-B2 CrossCodeEval with BM25 Retrieval Strategy
+As shown in Table III, SynthCoder-8B-DPO demonstrates a modest overall improvement compared to its base model, and generally achieves the best performance among all models tested.
 This includes models with a larger parameter count, such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B.
-Specifically, the fine-tuned SynthCoder-8B-SFT shows an overall increase of 1.1% in EM, 0.6% in ES, and 0.7% in ES_R.
+Specifically, the fine-tuned SynthCoder-8B-SFT shows an overall increase of 2.1% in EM, 0.9% in ES, and 0.6% in ES_R.
 While there was a slight decrease in the metrics for Java, all other languages showed improvement.
-The aligned SynthCoder-8B-DPO-wo-RS achieves gains of 0.6% to 0.7% overall, and also secures the highest scores on Java within the SynthCoder-8B series.
-The final version with repetition suppression, SynthCoder-8B-DPO-RS, obtains slight improvements of 1.5%, 1.5%, and 1.1% in overall EM, ES, and ES_R, respectively, when compared to the base model.
+The aligned SynthCoder-8B-DPO-wo-RS achieves gains of 0.6% to 0.7% overall, and also secures the highest scores on Java within the SynthCoderseries.
+The final version with repetition suppression, SynthCoder-8B-DPO, obtains slight improvements of 1.5%, 1.5%, and 1.1% in overall EM, ES, and ES_R, respectively, when compared to the base model Seed-Coder-8B-Base.
 TABLE III: Performance of models on CrossCodeEval with BM25 retrieval strategy [19].
@@
 Qwen2.5-Coder-7B 61.5 87.8 84.2 50.1 83.7 79.3 45.4 79.8 74.2 47.9 83.7 78.6 50.1 83.4 78.6
-Seed-Coder-8B-Inst 60.2 $\downarrow$8.9% 88.7 $\downarrow$2.8% 84.8 $\downarrow$3.3% 49.1 $\downarrow$5.0% 82.6 $\downarrow$3.0% 78.2 $\downarrow$2.8% 48.6 $\downarrow$1.9% 81.6 $\downarrow$1.5% 76.3 $\downarrow$1.1% 48.5 $\downarrow$6.5% 84.0 $\downarrow$2.8% 79.0 $\downarrow$2.7% 50.7 $\downarrow$5.5% 83.9 $\downarrow$2.5% 79.1 $\downarrow$2.4%
-Seed-Coder-8B-Base 61.9 $\downarrow$5.9% 89.1 $\downarrow$2.3% 85.4 $\downarrow$2.5% 52.6 $\uparrow$1.9% 84.5 $\downarrow$0.8% 80.4 -0.0% 49.3 $\downarrow$0.3% 81.9 $\downarrow$1.2% 76.8 $\downarrow$0.4% 50.6 $\downarrow$1.9% 84.8 $\downarrow$1.8% 80.0 $\downarrow$1.4% 52.7 $\downarrow$1.5% 84.7 $\downarrow$1.5% 80.2 $\downarrow$1.1%
-SynthCoder-8B-SFT 63.5 $\downarrow$3.3% 89.8 $\downarrow$1.4% 86.3 $\downarrow$1.4% 52.0 $\uparrow$0.8% 84.4 $\downarrow$0.9% 80.2 $\downarrow$0.2% 49.8 $\uparrow$0.6% 82.4 $\downarrow$0.6% 77.4 $\uparrow$0.3% 51.5 -0.0% 85.5 $\downarrow$1.0% 80.8 $\downarrow$0.4% 53.3 $\downarrow$0.4% 85.2 $\downarrow$0.9% 80.7 $\downarrow$0.4%
-SynthCoder-8B-DPO-wo-RS 65.4 $\downarrow$0.2% 91.1 -0.0% 87.5 $\downarrow$0.1% 52.9 $\uparrow$2.5% 85.4 $\uparrow$0.3% 80.8 $\uparrow$0.5% 48.7 $\downarrow$1.5% 82.4 $\downarrow$0.6% 76.6 $\downarrow$0.8% 51.7 $\uparrow$0.3% 86.1 $\downarrow$0.3% 80.9 $\downarrow$0.2% 53.6 $\uparrow$0.2% 85.8 $\downarrow$0.2% 80.9 $\downarrow$0.2%
-SynthCoder-8B-DPO-RS 65.5 91.1 87.6 51.6 85.1 80.4 49.5 82.9 77.2 51.6 86.3 81.1 53.5 86.0 81.0
-IV-B2 CoLT with Dependency-based Strategy Only
-The test set of CoLT-132K is a benchmark proposed in the aiXcoder-7B-v2 [10] to evaluate a model’s ability to utilize cross-file context for code completion.
-It encompasses Python, Java, C/C++, and Go with three completion scenarios: API calls, single-line completions, and span completions, comprising a total of 12,000 samples.
-In this work, we use Exact Match (EM) and Edit Similarity (ES) as the evaluation metrics.
-To specifically assess the model’s ability to leverage dependency-based cross-file context, we exclusively retain the context retrieved through dependency call relationships for each test case.
-For each model, we employ its corresponding prompt format for completion, with standardized total input context length of 4096 tokens, maximum output length of 200 tokens, and temperature of 0.
-As demonstrated in Table IV, SynthCoder-8B shows improvement over its base model in all four programming languages and achieves the best performance among all tested models.
+Seed-Coder-8B-Inst 60.2 $\uparrow$8.9% 88.7 $\uparrow$2.8% 84.8 $\uparrow$3.3% 49.1 $\uparrow$5.0% 82.6 $\uparrow$3.0% 78.2 $\uparrow$2.8% 48.6 $\uparrow$1.9% 81.6 $\uparrow$1.5% 76.3 $\uparrow$1.1% 48.5 $\uparrow$6.5% 84.0 $\uparrow$2.8% 79.0 $\uparrow$2.7% 50.7 $\uparrow$5.5% 83.9 $\uparrow$2.5% 79.1 $\uparrow$2.4%
+Seed-Coder-8B-Base 61.9 $\uparrow$5.9% 89.1 $\uparrow$2.3% 85.4 $\uparrow$2.5% 52.6 $\downarrow$1.9% 84.5 $\uparrow$0.8% 80.4 -0.0% 49.3 $\uparrow$0.3% 81.9 $\uparrow$1.2% 76.8 $\uparrow$0.4% 50.6 $\uparrow$1.9% 84.8 $\uparrow$1.8% 80.0 $\uparrow$1.4% 52.7 $\uparrow$1.5% 84.7 $\uparrow$1.5% 80.2 $\uparrow$1.1%
+SynthCoder-8B-SFT 63.5 $\uparrow$3.3% 89.8 $\uparrow$1.4% 86.3 $\uparrow$1.4% 52.0 $\downarrow$0.8% 84.4 $\uparrow$0.9% 80.2 $\uparrow$0.2% 49.8 $\downarrow$0.6% 82.4 $\uparrow$0.6% 77.4 $\downarrow$0.3% 51.5 -0.0% 85.5 $\uparrow$1.0% 80.8 $\uparrow$0.4% 53.3 $\uparrow$0.4% 85.2 $\uparrow$0.9% 80.7 $\uparrow$0.4%
+SynthCoder-8B-DPO-wo-RS 65.4 $\uparrow$0.2% 91.1 -0.0% 87.5 $\uparrow$0.1% 52.9 $\downarrow$2.5% 85.4 $\downarrow$0.3% 80.8 $\downarrow$0.5% 48.7 $\uparrow$1.5% 82.4 $\uparrow$0.6% 76.6 $\uparrow$0.8% 51.7 $\downarrow$0.3% 86.1 $\uparrow$0.3% 80.9 $\uparrow$0.2% 53.6 $\downarrow$0.2% 85.8 $\uparrow$0.2% 80.9 $\uparrow$0.2%
+SynthCoder-8B-DPO 65.5 91.1 87.6 51.6 85.1 80.4 49.5 82.9 77.2 51.6 86.3 81.1 53.5 86.0 81.0
+IV-B3 CoLT with Dependency-based Strategy Only
+As demonstrated in Table IV, SynthCoder-8B-DPO shows improvement over its base model in all four programming languages and achieves the best performance among all tested models.
 This includes outperforming models with larger parameter counts, such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B.
-The fine-tuned SynthCoder-8B-SFT improves the overall EM and ES by 4.1% and 4.2%, respectively, with gains across all individual metrics, including a maximum improvement of 7.4% on the Go EM score.
-The aligned SynthCoder-8B-DPO-wo-RS further enhances performance over SynthCoder-8B-SFT on all metrics, with an additional increase of 1.5% in overall EM and 1.4% in overall ES.
-The final version with repetition suppression, SynthCoder-8B-DPO-RS, achieves the optimal results across all metrics, ultimately boasting a 7.0% improvement in overall EM and a 7.5% improvement in overall ES compared to the base model.
-It is worth noting that for all metrics on this benchmark, each subsequent stage of our process – fine-tuning, alignment, and repetition suppression – resulted in a consistent improvement in the performance of the SynthCoder-8B model.
+The fine-tuned SynthCoder-8B-SFT improves the overall EM and ES by 3.9% and 3.9%, respectively, with gains across all individual metrics, including a maximum improvement of 7.3% on the Go EM score.
+The aligned SynthCoder-8B-DPO-wo-RS further enhances performance over SynthCoder-8B-SFT on all metrics, with an additional increase of 1.5% in overall EM and 1.0% in overall ES.
+The final version with repetition suppression, SynthCoder-8B-DPO, achieves the optimal results across all metrics, ultimately boasting a 7.0% improvement in overall EM and a 7.5% improvement in overall ES compared to the base model.
+It is worth noting that for all metrics on this benchmark, each subsequent stage of our process resulted in a consistent improvement in the final.
 TABLE IV: Performance of models on CoLT with dependency-based strategy only [10].
@@
 Qwen2.5-Coder-7B 35.8 58.5 47.6 74.7 47.7 65.1 38.5 66.5 42.4 66.2
-Seed-Coder-8B-Inst 42.8 $\downarrow$16.0% 71.7 $\downarrow$12.1% 48.5 $\downarrow$15.2% 76.2 $\downarrow$9.9% 57.7 $\downarrow$8.2% 81.1 $\downarrow$6.3% 46.0 $\downarrow$12.5% 75.1 $\downarrow$8.3% 48.8 $\downarrow$12.7% 76.0 $\downarrow$9.1%
-Seed-Coder-8B-Base 46.4 $\downarrow$7.0% 74.2 $\downarrow$8.3% 51.0 $\downarrow$9.7% 76.6 $\downarrow$9.3% 59.6 $\downarrow$4.6% 81.6 $\downarrow$5.7% 48.4 $\downarrow$6.9% 76.3 $\downarrow$6.6% 51.4 $\downarrow$7.0% 77.2 $\downarrow$7.5%
-SynthCoder-8B-SFT 47.5 $\downarrow$4.5% 76.6 $\downarrow$4.9% 54.7 $\downarrow$2.3% 81.3 $\downarrow$3.1% 60.8 $\downarrow$2.6% 83.8 $\downarrow$2.9% 50.5 $\downarrow$2.4% 79.4 $\downarrow$2.5% 53.4 $\downarrow$2.9% 80.2 $\downarrow$3.3%
-SynthCoder-8B-DPO-wo-RS 48.9 $\downarrow$1.6% 78.6 $\downarrow$2.2% 55.1 $\downarrow$1.6% 82.0 $\downarrow$2.1% 62.1 $\downarrow$0.5% 85.3 $\downarrow$1.1% 50.6 $\downarrow$2.2% 79.6 $\downarrow$2.2% 54.2 $\downarrow$1.4% 81.4 $\downarrow$1.9%
-SynthCoder-8B-DPO-RS 49.7 80.3 55.9 83.8 62.4 86.2 51.7 81.4 54.9 82.9
-IV-B3 ExecRepoBench
-The ExecRepoBench benchmark [^31] https://huggingface.co/datasets/CSJianYang/ExecRepoBench, proposed by the Alibaba team, focuses on evaluating the code completion capabilities of LLMs in real-world software development environments.
-The dataset is sourced from 50 actively maintained Python repositories on GitHub, comprising a total of 1,200 test cases.
-For test case construction, the benchmark introduces a multi-level syntax masking method based on Abstract Syntax Trees (AST).
-It utilizes Tree-sitter to classify syntax nodes into three categories – functions, statements, and expressions – to create test instances.
-Additionally, it employs non-syntactic construction methods, including snippet, single-line, and multi-line truncation.
-The benchmark provides not only the intra-file prefix and suffix in a Fill-in-the-Middle (FIM) format but also the relevant cross-file context from the same repository in the form of complete files.
-In our reproduction of the ExecRepoEval evaluation, we implemented a rule-based fix to correct indentation errors in the code generated by the models.
-We use Pass@1 and Edit Similarity (ES) to measure model performance.
-For each model, we employed its corresponding prompt format for completion, with limits of 4096 tokens for the intra-file context, 4096 tokens for the cross-file context, and a maximum output length of 160 tokens.
-The temperature was set to 0.
-As shown in Table V, SynthCoder-8B achieves a 24.5% improvement in Pass@1 and a 35.3% improvement in ES compared to its base model.
-Its performance is also significantly higher than that of other 7B and 14B models, including Qwen2.5-Coder-14B-Inst.
+Seed-Coder-8B-Inst 42.8 $\uparrow$16.0% 71.7 $\uparrow$12.1% 48.5 $\uparrow$15.2% 76.2 $\uparrow$9.9% 57.7 $\uparrow$8.2% 81.1 $\uparrow$6.3% 46.0 $\uparrow$12.5% 75.1 $\uparrow$8.3% 48.8 $\uparrow$12.7% 76.0 $\uparrow$9.1%
+Seed-Coder-8B-Base 46.4 $\uparrow$7.0% 74.2 $\uparrow$8.3% 51.0 $\uparrow$9.7% 76.6 $\uparrow$9.3% 59.6 $\uparrow$4.6% 81.6 $\uparrow$5.7% 48.4 $\uparrow$6.9% 76.3 $\uparrow$6.6% 51.4 $\uparrow$7.0% 77.2 $\uparrow$7.5%
+SynthCoder-8B-SFT 47.5 $\uparrow$4.5% 76.6 $\uparrow$4.9% 54.7 $\uparrow$2.3% 81.3 $\uparrow$3.1% 60.8 $\uparrow$2.6% 83.8 $\uparrow$2.9% 50.5 $\uparrow$2.4% 79.4 $\uparrow$2.5% 53.4 $\uparrow$2.9% 80.2 $\uparrow$3.3%
+SynthCoder-8B-DPO-wo-RS 48.9 $\uparrow$1.6% 78.6 $\uparrow$2.2% 55.1 $\uparrow$1.6% 82.0 $\uparrow$2.1% 62.1 $\uparrow$0.5% 85.3 $\uparrow$1.1% 50.6 $\uparrow$2.2% 79.6 $\uparrow$2.2% 54.2 $\uparrow$1.4% 81.4 $\uparrow$1.9%
+SynthCoder-8B-DPO 49.7 80.3 55.9 83.8 62.4 86.2 51.7 81.4 54.9 82.9
+IV-B4 ExecRepoBench
+As shown in Table V, SynthCoder-8B-DPO achieves significant improvement in Pass@1 and ES compared to its base model, higher than that of other 7B models.
 We did not include Qwen2.5-Coder-Instruct-C, which was proposed in the original ExecRepoBench paper, as its weights have not been publicly released.
-The fine-tuned SynthCoder-8B-SFT improves upon the base model by 8.4% in Pass@1 and 8.6% in ES.
-The aligned SynthCoder-8B-DPO-wo-RS further boosts performance by 15.5% in Pass@1 and 25.1% in ES relative to SynthCoder-8B-SFT.
-The final version with repetition suppression, SynthCoder-8B-DPO-RS, adds an additional 0.6% to Pass@1 and 1.6% to ES, culminating in a total improvement of 24.5% and 35.3% over the original base model.
+The fine-tuned SynthCoder-8B-SFT improves upon the base model by 7.2% in Pass@1 and 6.8% percentage points in ES.
+The aligned SynthCoder-8B-DPO-wo-RS further boosts performance by 6.9 percentage points in Pass@1 and 12.5 percentage points in ES relative to SynthCoder-8B-SFT.
+The final version with repetition suppression, SynthCoder-8B-DPO, adds an additional 0.6% to Pass@1 and 1.6% to ES, culminating in a total improvement of 24.5% and 35.3% over Seed-Coder-8B-Base.
 TABLE V: Performance of models on ExecRepoEval [12].
@@
 Qwen2.5-Coder-7B 18.2 19.4
-Seed-Coder-8B-Inst 32.8 $\downarrow$89.3% 38.1 $\downarrow$36.0%
-Seed-Coder-8B-Base 45.9 $\downarrow$35.3% 41.6 $\downarrow$24.5%
-SynthCoder-8B-SFT 49.0 $\downarrow$26.7% 44.6 $\downarrow$16.1%
-SynthCoder-8B-DPO-wo-RS 61.1 $\downarrow$1.6% 51.5 $\downarrow$0.6%
-SynthCoder-8B-DPO-RS 62.1 51.8
-IV-C Errors of Wrongly Repeating Suffix or Prefix
+Seed-Coder-8B-Inst 32.8 $\uparrow$89.3% 38.1 $\uparrow$36.0%
+Seed-Coder-8B-Base 45.9 $\uparrow$35.3% 41.6 $\uparrow$24.5%
+SynthCoder-8B-SFT 49.0 $\uparrow$26.7% 44.6 $\uparrow$16.1%
+SynthCoder-8B-DPO-wo-RS 61.1 $\uparrow$1.6% 51.5 $\uparrow$0.6%
+SynthCoder-8B-DPO 62.1 51.8
+IV-C Tendency to Repeat Context
 When applying base models to practical code completion plugins, we observed that models frequently tend to erroneously repeat the preceding (prefix) or succeeding (suffix) context.
 This behavior can frustrate users and lead them to question the capabilities of both the model and the plugin.
-To analyze this issue systematically, we focus on a specific scenario: cases where the first non-empty line of the generated content does not exactly match the expected completion, but is identical to the last line of the prefix or the first line of the suffix after whitespace is removed.
+To analyze this issue systematically, we focus on a specific scenario: cases where the first non-empty line of the generated content does not exactly match the expected completion’s first non-empty line, but is identical to the last non-empty line of the prefix or the first non-empty line of the suffix after whitespace is removed.
 Our models significantly reduce the probability of such anomalous repetitions.
@@
 Qwen2.5-Coder-7B 0.69% 0.09% 1.51%
-Seed-Coder-8B-Inst 2.74% $\uparrow$621% 0.43% $\uparrow$377% 2.54% $\uparrow$418%
-Seed-Coder-8B-Base 1.67% $\uparrow$339% 2.23% $\uparrow$2377% 1.77% $\uparrow$261%
-SynthCoder-SFT 0.60% $\uparrow$57% 2.58% $\uparrow$2766% 1.42% $\uparrow$189%
-SynthCoder-DPO-wo-RS 0.52% $\uparrow$36% 0.60% $\uparrow$566% 0.50% $\uparrow$2%
-SynthCoder-DPO-RS 0.38% 0.09% 0.49%
+Seed-Coder-8B-Inst 2.74% 0.43% 2.54%
+Seed-Coder-8B-Base 1.67% 2.23% 1.77%
+SynthCoder-SFT 0.60% 2.58% 1.42%
+SynthCoder-DPO-wo-RS 0.52% 0.60% 0.50%
+SynthCoder-8B-DPO 0.38% 0.09% 0.49%
     AX: aiXcoder; ERE: ExecRepoEval.
@@
 Qwen2.5-Coder-7B 0.18% 0.43% 0.04%
-Seed-Coder-8B-Inst 0.33% $\uparrow$560% 0.09% -0% 0.08% $\uparrow$300%
-Seed-Coder-8B-Base 0.56% $\uparrow$1020% 0.09% -0% 0.08% $\uparrow$300%
-SynthCoder-SFT 0.15% $\uparrow$200% 0.34% $\uparrow$278% 0.16% $\uparrow$700%
-SynthCoder-DPO-wo-RS 0.10% $\uparrow$100% 0.17% $\uparrow$89% 0.08% $\uparrow$300%
-SynthCoder-DPO-RS 0.05% 0.09% 0.02%
+Seed-Coder-8B-Inst 0.33% 0.09% 0.08%
+Seed-Coder-8B-Base 0.56% 0.09% 0.08%
+SynthCoder-SFT 0.15% 0.34% 0.16%
+SynthCoder-DPO-wo-RS 0.10% 0.17% 0.08%
+SynthCoder-8B-DPO 0.05% 0.09% 0.02%
     AX: aiXcoder; ERE: ExecRepoEval.
-We analyzed the frequency of erroneous suffix and prefix repetitions in the generated outputs across the aforementioned datasets, with the results presented in Table VI and Table VII, respectively.
-The CrossCodeEval dataset is omitted from these tables because its ground-truth completions do not contain full lines, meaning the specific type of whole-line repetition we define does not occur.
-We observed that for models using either the Prefix-Suffix-Middle (PSM) FIM format (e.g., the Qwen-Coder and DeepSeek series) or the Suffix-Prefix-Middle (SPM) FIM format (e.g., the Seed-Coder series), the frequency of erroneously repeating the suffix is generally higher than that of repeating the prefix.
-The StarCoder2-7B model exhibits a particularly high repetition rate, reaching a suffix repetition rate of 30.58% on the ExecRepoEval benchmark.
-While the prefix repetition rate of DeepSeek-Coder is generally lower than that of the Qwen-Coder series, the opposite is true for their suffix repetition rates.
-Noteworthly, The Seed-Coder series of models demonstrates lower repetition rates compared to the other model families.
+We analyzed the frequency of erroneous suffix and prefix repetitions across the evaluated datasets, with detailed results presented in Table VI and Table VII.
+The CrossCodeEval dataset was excluded from this analysis because its completion postitions are always in the middle of a line, which precludes the specific type of whole-line repetition under investigation.
+Particularly pronounced for StarCoder2-7B, which reached a suffix repetition rate of 30.58% on the ExecRepoEval benchmark.
+Furthermore, while the DeepSeek-Coder series demonstrates a lower prefix repetition rate than the Qwen-Coder series, its suffix repetition rate is conversely higher.
 We also noted that comparing the model before and after fine-tuning (i.e., Seed-Coder-8B-Base vs.
@@
 This is because outputs containing repetitions that are generated and rejected during the sampling process are subsequently used as the negative examples in the preference pairs for DPO training.
-After incorporating an additional 11% of repetition-suppression data, SynthCoder-DPO-RS demonstrated a significantly greater reduction in this behavior.
+After incorporating an additional 11% of repetition-suppression data, SynthCoder-8B-DPO demonstrated a significantly greater reduction in this behavior.
 It achieved a prefix repetition rate below 0.5% and a suffix repetition rate below 0.1% across all datasets, marking the best performance among the evaluated models.
@@
 Addressing these challenges requires a concerted effort, spanning from the curation of training data on the model side to the implementation of output filtering mechanisms on the tool side.
-We also observed significant variations in output length among different models given the same input.
-Ideally, generated code should be concise and to the point, but this is often not the case.
-Models without fine-tuning are prone to generating excessively long outputs, either by endlessly repeating permutations of the surrounding code or by arbitrarily adding extensive, non-essential docstrings.
-This can lead to the model reaching its maximum generation length, resulting in unusable code.
-This behavior is particularly pronounced in the Qwen-Coder series of models and is frequently observed in the ExecRepoEval benchmark.
-For developers, such unbounded generation not only needlessly consumes computational resources on the model side but also wastes their time and mental energy waiting for and evaluating the results.
-This underscores the importance of constructing training corpora based on syntax tree nodes and user behavior patterns, which enables the model to generate more targeted and appropriately sized completions.
-These crucial capabilities are not adequately captured by current evaluation benchmarks.
 In summary, enhancing the performance of code completion necessitates a collaborative effort between LLM and client.
@@
 Moreover, we maintain consistency between the data filtering methods used during training and those applied during inference, which enhances the project-level code completion capabilities of SynthCoder in practical application scenarios.
-V-C Next Edit Suggestions
-Next Edit Prediction infers a developer’s intent by analyzing the code context and historical edit sequences.
-It can predict the content and location of the next edit without requiring natural language input, thereby preventing interruptions to the workflow [45].
-For instance, during repetitive tasks such as API updates or identifier renaming, it can proactively suggest coherent subsequent edits based on the user’s initial modification.
-This feature has already been implemented in GitHub Copilot [^32] https://code.visualstudio.com/blogs/2025/02/12/next-edit-suggestions, where it works in synergy with traditional code completion capabilities.
-The core of the edit prediction task is to modify existing code: whether to fix bugs or iterate on features, rather than to rewrite it entirely.
-This area has been a subject of extensive academic research.
-For example, CoditT5 [46] introduced a pre-trained model specifically designed for code editing.
-Modit [47] developed a multi-modal editing engine that integrates edit location, code context, and commit information.
-More recent approaches like jLED [48] employ a joint training methodology, enabling the model to simultaneously learn where to edit and how to edit, providing a more practical solution for real-world editing scenarios.
-Collectively, these research efforts have advanced the ability of models to follow precise instructions and execute complex code modifications.
-In contrast to the fine-grained task of performing insertions, deletions, and modifications at the cursor, many current methods for evaluating the coding capabilities of large models focus on code generation, which does not fully capture the nuances of edit prediction.
-For example, mainstream function-level benchmarks, including HumanEval [49] and MBPP [50], require models to generate complete code blocks based on function signatures, docstrings, or functional descriptions.
-The limitation of such benchmarks is that they measure a model’s ability to comprehend high-level instructions and generate code from scratch, rather than its capacity for making precise, intelligent, and incremental modifications within the complex context of existing code.
-Meanwhile, existing code completion datasets primarily focus on token-level completion at the cursor based on the surrounding code context; they do not incorporate edit history, making it difficult to assess a model’s ability to reason by analogy from past edits.
-Looking ahead, models such as zeta [^33] https://huggingface.co/zed-industries/zeta have already successfully integrated FIM-based code completion with edit prediction, achieving consistently strong performance across both capabilities.
-This integration will enable models to become more intelligent and better aligned with developer workflows, driving a paradigm shift in code completion from a reactive response to a proactive prediction model that incorporates edit suggestions.
 VI Threats to Validity
 In this section, we discuss some factors that may potentially bring risks to the findings of the study.
-Limited models
-Actually, we established SynthCoder based on only Seed-Coder-8B-Base as it is the current state-of-the-art (SOTA) LLM in code completion so far.
-However, SynthCoder is designed as a model-agnostic approach, therefore any LLM with capability of conducting FIM code completion tasks can be post-trained to create variations of SynthCoder.
-In this sense, it raises the necessity to replicate our approaches using other LLM series.
-Limited datasets
-The evaluation of SynthCoder is based on a limited number of datasets and benchmark methods, which may bring in certain validity risks.
-In particular, the test sets may not fully cover the diversity of real-world application scenarios, leading to potential gaps between benchmark results and practical performance.
-However, we have alreadly selected industry-recognized open-source benchmarks (ExecRepoEval and CrossCodeEval) and a custom test set constructed from third-party open-source datasets.
-Meanwhile, we evaluated the capabilities of SynthCoder and other baseline models in both single-file code completion, cross-file code completion based on dependency graph retrieval, and cross-file code completion based on BM25 retrieval so as to establish a comprehensive understanding of the performance of SynthCoderṪhe results strongly demonstrate the superiority of our method and model.
-Throughout the evaluation process, we utilized the official testing frameworks and parameters provided by the benchmarks or other well-known open-source models wherever possible to ensure the validity and generalizability of our tests.
 Data Leakage
@@
 By constructing and comparing a repository-file tuple structure, we controlled the possibility of data contamination resulting from overlapping files between the training and test sets.
+Training Hyperparameter
+Hyperparameters can significantly impact model performance.
+Due to computational constraints, we used the officially recommended settings for each benchmark model during comparisons rather than aligning hyperparameters across all methods.
+In this sense, we cannot entirely rule out the possibility that the performance gains we observed were solely attributable to the dataset we constructed.
+Nevertheless, we believe that the hyperparameter configurations used by other benchmark methods likely represented their respective optimal setups at the time.
+We encourage interested researchers to reproduce our work and explore potentially better hyperparameter configurations.
+Limited Models
+Actually, we established SynthCoder based on only Seed-Coder-8B-Base as it is the current state-of-the-art (SOTA) LLM in code completion so far.
+However, SynthCoder is designed as a model-agnostic approach, therefore any LLM with capability of conducting FIM code completion tasks can be post-trained to create variations of SynthCoder.
+In this sense, it raises the necessity to replicate our approaches using other LLM series.
+Limited Benchmark Datasets
+The evaluation of SynthCoder is based on a limited number of datasets and benchmark methods, which may bring in certain validity risks.
+In particular, the test sets may not fully cover the diversity of real-world application scenarios, leading to potential gaps between benchmark results and practical performance.
+However, we have alreadly selected industry-recognized open-source benchmarks.
+Meanwhile, we evaluated the capabilities of SynthCoder and other baseline models in both single-file code completion, cross-file code completion based on dependency graph retrieval, and cross-file code completion based on BM25 retrieval so as to establish a comprehensive understanding of the performance of SynthCoder .
+The results strongly demonstrate the superiority of our method and model.
+Throughout the evaluation process, we utilized the well-known testing frameworks and official parameters wherever possible to ensure the validity and generalizability of our tests.
 VII Conclusion
@@
 Furthermore, we discuss optimization directions for deploying such models in production environments that extend beyond conventional evaluation metrics.
+Building on our findings, we identify several promising directions for future research.
+Future work will advance beyond FIM towards more sophisticated edit prediction by learning from real-world code changes [50].
+The practical deployment of such models hinges on a synergistic approach: creating compact, specialized models via techniques like knowledge distillation [51], and coupling them with system-level optimizations.
+This includes leveraging speculative decoding for low latency [52], paving the way for faster, smarter, and more dependable code assistants.